### v1

# Lattice isomorphic Banach lattices of polynomials

Abstract
We study Díaz-Dineen’s problem for regular homogeneous vector-valued polynomials. In particular, we prove that whenever $E^{*}$ and $F^{*}$ are lattice isomorphic with at least one having order continuous norm, then $\mathcal{P}^{r}(^{n}E;G^{*})$ and $\mathcal{P}^{r}(^{n}F;G^{*})$ are lattice isomorphic for every $n\in\mathbb{N}$ and every Banach lattice $G$. We also study the analogous problem for the classes of regular compact, regular weakly compact, orthogonally additive and regular nuclear polynomials.
1 Introduction and background
In [15], Díaz and Dineen posed the following question: If $X$ and $Y$ are dual isomorphic Banach spaces, that is $X^{\ast}$ and $Y^{\ast}$ are isomorphic as Banach spaces, are $\mathcal{P}(^{n}X)$ and $\mathcal{P}(^{n}Y)$ isomorphic as Banach spaces? Although it is not known if the above question has a positive answer, some partial answers were presented in [13, 15, 22]. In particular, J. C. Díaz and S. Dineen proved that the problem has a positive solution whenever $X^{\ast}$ has both the Schur property and the approximation property (see [15, Proposition 4]). In [13], F. Cabello Sánchez, J. Castillo and R. García generalized Díaz-Dineen results by proving a positive solution for this problem whenever $X$ is Arens-regular or whenever $X$ and $Y$ are stable. The same result was proven, independently, by S. Lassalle and I. Zalduendo in [22] who also addressed the same question concerning other classes polynomials. Further, in [14], D. Carando and S. Lassalle extended these results to spaces of vector-valued homogeneous polynomials.
Motivated by the increasing interest in studying homogeneous polynomials on Banach lattices (see, e.g., [5, 6, 7, 8, 9, 10, 18, 24]), the following question arises naturally: Given two dual isomorphic Banach lattices $E$ and $F$, that is $E^{\ast}$ and $F^{\ast}$ are lattice isomorphic, under which conditions $\mathcal{P}^{r}(^{n}E)$ and $\mathcal{P}^{r}(^{n}F)$ are lattice isomorphic for every $n\in\mathbb{N}$? By adapting the proofs from [14] and [22], we were able to provide a positive answer whenever $E^{*}$ and $F^{*}$ have order continuous norms. In particular, we prove that whenever $E^{\ast}$ and $F^{\ast}$ are lattice isomorphic with at least one (and hence the other) having order continuous norm then $\mathcal{P}^{r}(^{n}E;G^{*})$ and $\mathcal{P}^{r}(^{n}F;G^{*})$ are lattice isomorphic for every $n\in\mathbb{N}$ and every Banach lattice $G$. We observe that, in our setting, the assumption of Arens regularity is replaced with the requirement that the norm on the dual is order continuous. This replacement is justified by the fact that a positive version of Arens regularity for $E$ is equivalent to the order continuity of the norm on $E^{*}$ having order continuous norm. Furthermore, as we will demonstrate in Proposition 2.1, positive Arens regularity coincides with positive symmetric Arens regularity, highlighting a noteworthy contrast with the Banach space case.
As an application, we obtain conditions so that $\mathcal{P}_{\cal K}^{r}(^{n}E;G^{*})$ (resp. $\mathcal{P}_{\cal W}^{r}(^{n}E;G^{*})$) and $\mathcal{P}_{\cal K}^{r}(^{n}F;G^{*})$ (resp. $\mathcal{P}_{\cal W}^{r}(^{n}F;G^{*})$) are isomorphic, where $\mathcal{P}_{\cal K}^{r}$ and $\mathcal{P}_{\cal W}^{r}$ denote, respectively, the classes of regular compact and regular weakly compact polynomials. We also consider the class of orthogonally additive polynomials. In Section 3, we investigate this problem for a new class of polynomials. Specifically, inspired by ideas from [4], we define the class of regular nuclear $n$-homogeneous polynomials, denoted by $\mathcal{P}_{\cal N}^{r}$, and prove that whenever $G$ is Dedekind complete with the lattice approximation property and $E^{*}$ is lattice isomorphic to $F^{*}$, then $\mathcal{P}_{\cal N}^{r}(^{n}E;G)$ and $\mathcal{P}_{\cal N}^{r}(^{n}F;G)$ are lattice isomorphic for every $n\in\mathbb{N}$.
Before proceeding, we introduce some definitions and notation. Throughout this paper, $X$ and $Y$ denote Banach spaces, $E$ and $F$ denote Banach lattices, $E^{+}$, $B_{E}$ and $S_{E}$ denote, respectively, the positive cone, the closed unit ball, and the unit sphere of $E$. Given Banach spaces $X_{1},\dots,X_{n}$ and $Y$, the Banach space of all continuous $n$-linear operators $A\colon X_{1}\times\cdots\times X_{n}\to Y$ is denoted by $\mathcal{L}(X_{1},\dots,X_{n};Y)$. In the case that $F=\mathbb{K}$, we denote simply by $\mathcal{L}(X_{1},\dots,X_{n})$. For given Banach lattices $E_{1},\dots,E_{n}$ and $F$, a $n$-linear operator $A\colon E_{1}\times\cdots\times E_{n}\to F$ is said to be positive if $A(x_{1},\dots,x_{n})\geq 0$ for all $x_{1}\in E_{1}^{+},\dots,x_{n}\in E_{n}^{+}$. The difference of two positive $n$-linear operators is called a regular $n$-linear operator, and the set of all regular $n$-linear operators from $E_{1}\times\cdots\times E_{n}$ into $F$ is denoted by $\mathcal{L}^{r}(E_{1},\dots,E_{n};F)$. Whenever $F$ is Dedekind complete, $\mathcal{L}^{r}(E_{1},\dots,E_{n};F)$ is a Banach lattice with the regular norm $\|A\|_{r}=\||A|\|$, where $|A|$ denotes the absolute value of the regular $n$-linear operator $A\colon E_{1}\times\cdots\times E_{n}\to F$. The positive projective tensor product, or the Fremlin tensor product of $E_{1},\dots,E_{n}$, denoted by $E_{1}\widehat{\otimes}_{|\pi|}\cdots\widehat{\otimes}_{|\pi|}E_{n}$, is the Banach lattice generated by considering the completion of the vector lattice tensor product $E_{1}\overline{\otimes}\cdots\overline{\otimes}\,E_{n}$ with respect the positive projective tensor norm:
$\|u\|_{|\pi|}=\inf\left\{{\sum_{i=1}^{k}\|x_{i}^{1}\|\cdots\|x_{i}^{n}\|}\,:\,{|u|\leq\sum_{i=1}^{n}x_{i}^{1}\otimes\cdots\otimes x_{i}^{n},\,x_{i}^{j}\in E_{j}^{+}}\right\}.$
For each $T\in\mathcal{L}^{r}(E_{1},\dots,E_{m};F)$, there exists a unique bounded linear operator $T^{\otimes}$ from $E_{1}\widehat{\otimes}_{|\pi|}\cdots\widehat{\otimes}_{|\pi|}E_{n}$ into $F$, called the linearization of $T$, such that $T^{\otimes}(x_{1}\otimes\cdots\otimes x_{n})=T(x_{1},\dots,x_{n})$ for all $x_{1}\in E_{1},\dots,x_{n}\in E_{n}$. Moreover, the map
$T\in\mathcal{L}^{r}(E_{1},\dots,E_{m};F)\mapsto T^{\otimes}\in\mathcal{L}^{r}(E_{1}\widehat{\otimes}_{|\pi|}\cdots\widehat{\otimes}_{|\pi|}E_{n};F)$
is an isometric isomorphism and a lattice hommomorphism (see [11, Proposition 3.3]).
An $n$-homogeneous polynomial between Riesz spaces $P\colon E\to F$ is positive if its associated symmetric multilinear operator $T_{P}\colon E^{n}\to F$ is positive. The difference of two positive $n$-homogeneous polynomials is called a regular homogeneous polynomial, and the set of all these polynomials is denoted by $\mathcal{P}^{r}(^{n}E,F)$. When $F$ is the scalar field we simply write $\mathcal{P}^{r}(^{n}E)$. If $E$ and $F$ are Banach lattices with $F$ Dedekind complete, then $\mathcal{P}^{r}(^{n}E,F)$ is a Banach lattice with the regular norm $\|P\|_{r}=\||P|\|$, where $|P|$ denotes the absolute value of the regular $n$-homogeneous polynomial $P\colon E\to F$. The $n$-fold positive projective symmetric tensor product of $E$ by $\widehat{\otimes}_{n,s,|\pi|}E$, which is a Banach lattice endowed with the positive projective symmetric tensor norm:
$\|u\|_{s,|\pi|}=\inf\left\{{\sum_{i=1}^{k}\|x_{i}\|^{n}}\,:\,{|u|\leq\sum_{i=1}^{k}\otimes^{n}x_{i},\,x_{i}\in E^{+}}\right\},$
where $\otimes^{n}x=x\,\otimes\stackrel{{\scriptstyle n}}{{\cdots}}\otimes\,x$ for every $x\in E$. Let $\theta_{n}\colon E\to\widehat{\otimes}_{n,s,|\pi|}E$ be the canonical $n$-homogeneous polynomial given by $\theta_{n}(x)=\otimes^{n}x$. We note that $\theta_{n}$ is a lattice homomorphism. For every $P\in\mathcal{P}^{r}(^{n}E,F)$ there exists a unique regular linear operator $P^{\otimes}\colon\widehat{\otimes}_{n,s,|\pi|}E\to F$, called the linearization of $P$, such that $P(x)=P^{\otimes}(\otimes^{n}x)$ for every $x\in E$. The operator
$\Phi\colon\mathcal{L}^{r}\left(\widehat{\otimes}_{n,s,|\pi|}E;F\right)\to\mathcal{P}^{r}(^{n}E;F)~,~\Phi(T)=T\circ\theta_{n},\quad T\in\mathcal{L}^{r}(\widehat{\otimes}_{n,s,|\pi|}E;F),$
is an isometric isomorphism and a lattice homomorphism. For each $P\in\mathcal{P}^{r}(^{n}E;F)$, the regular linear operator $\Phi^{-1}(P)\colon\widehat{\otimes}_{n,s,|\pi|}E\to F$ is called the linearization of $P$, and it is denoted by $P^{\otimes}$ (see [11, Proposition 3.4]).
We refer the reader to [1, 26] for background on Banach lattices, to [17] for Banach space theory, to [16] for polynomials on Banach spaces, and to [11, 25] for positive/regular multilinear operators or polynomials.
2 Díaz-Dineen problem for regular polynomials
We begin this section recalling that the first Arens adjoint of $A\in\mathcal{L}(X_{1},\cdots,X_{n};Y)$ is the $n$-linear map $A^{*}\colon F^{*}\times E_{1}\times\cdots\times E_{n-1}\to E_{n}^{*}$ defined by
$A^{*}(y^{*},x_{1},\dots,x_{n-1})(x_{n})=y^{*}(A(x_{1},\dots,x_{n})).$
The second Arens adjoint $A^{**}$ of $A$ is the first Arens adjoint of $A^{*}$, that is the $n$-linear map $A^{**}\colon E_{n}^{**}\times F^{*}\times E_{1}\times\cdots\times E_{n-2}\to E_{n-1}^{*}$ given by
$A^{**}(x_{n}^{**},y^{*},x_{1},\dots,x_{n-1})(x_{n}^{*})=x_{n}^{**}(A^{*}(y^{*},x_{1},\dots,x_{n-1})).$
This process is repeated until we obtain $\overline{A}:=A^{(n+1)*}\colon E_{1}^{**}\times\cdots\times E_{n}^{**}\to F^{**}$, the $(n+1)$-Arens adjoint of $A$ which is called the Arens-extension of $A$ (see [8, Section 2] for more details). For the remainder of the paper, we shall call this extension the Arens extension, bearing in mind that it coincides with its Aron-Berner extension (see [3]). Whenever $X=Y$, the transpose of $A$ is the bilinear map $A^{t}\colon X\times X\to Z$ given by $A^{t}(x,y)=A(y,x)$ for all $x,y\in X$. Still under the assumption that $X=Y$, AA is said to be Arens-regular (or simply regular) if $A^{t***t}=A^{***}$. It is important to note that we have two different concepts of regularity for multilinear operators, one in the lattice sense presented in the Introduction and the one in the extension sense. It follows from [29, Theorem 2.2] that a bilinear form $A\colon X\times X\to\mathbb{R}$ is Arens-regular if and only if the bounded linear operator $S_{A}\colon X\to X^{*}$ given by $S_{A}(x)(y)=A(x,y)$ is weakly compact.
Recall that a Banach space $X$ is called Arens-regular if all linear operators $X\to X^{\ast}$ are weakly compact. To our knowledge, the first appearance of a positive version of Arens-regularity was in [12], when G. Buskes and R. Page gave necessary and sufficient conditions so that every positive bilinear form $E\times E\to\mathbb{R}$ is Arens-regular. The next result summarizes what we have about this notion.
Proposition 2.1.
For a Banach lattice $E$, the following are equivalent:
(1) $E^{\ast}$ has order continuous norm.
(2) $E$ does not contain a sublattice isomorphic to $\ell_{1}$.
(3) Every positive bilinear map $A\colon E\times E\to\mathbb{R}$ is Arens-regular.
(4) Every positive operator $T\colon E\to E^{*}$ is weakly compact.
(5) Every positive symmetric bilinear map $A\colon E\times E\to\mathbb{R}$ is Arens-regular.
Proof.
[…]
$A\in\mathcal{L}^{+}(E,E;\mathbb{R})\mapsto S_{A}\in\mathcal{L}^{+}(E;E^{\ast}).$
[…]
$B(x,y)=\sum_{j\text{ even}}\sum_{k=1}^{j-1}x_{j}y_{k}\,+\sum_{k\text{ even}}\sum_{j=1}^{k-1}x_{j}y_{k},\quad x=(x_{i})_{i},y=(y_{i})_{i}\in\ell_{1}.$
Since $B$ is positive, we may consider the positive symmetric bilinear map $A\colon E\times E\to\mathbb{R}$ given by $A(x,y)=B(\Pi(x),\Pi(y))$ for all $x,y\in E$. Let us see that $A$ cannot be Arens-regular, showing that its associated operator $S_{A}\colon E\to E^{*}$ is not weakly compact: as $\{S_{B}(e_{2n}):n\in\mathbb{N}\}$ is not a weakly precompact subset of $\ell_{1}^{*}=\ell_{\infty}$ (see [3, p. 83]), there exists $\varphi\in\ell_{\infty}$ such that $(\varphi(S_{B}(e_{2n})))_{n}$ does not have any Cauchy subsequence. If $x_{n}\in E$ is such that $\Pi(x_{n})=e_{2n}$, then $(\varphi(S_{B}(\Pi(x_{n}))))_{n}$ does not have any Cauchy subsequence. By extending $\varphi\colon\ell_{1}\to\mathbb{R}$ to $E$ in such a way that it vanishes on the complement, we obtain that $\{\Pi^{*}\circ S_{B}\circ\Pi(x_{n}):n\in\mathbb{N}\}$ is not weakly precompact in $E^{*}$, and since $S_{A}=\Pi^{*}\circ S_{B}\circ\Pi$, we conclude that $S_{A}$ is not weakly compact. ∎
The fourth condition in Proposition 2.1 induces a positive version of Arens-regularity. However, since this is equivalent to a Banach lattice whose dual has order continuous norm, we will not introduce a new terminology. It is noteworthy mentioning that item (5) in Proposition 2.1 shows that the notions of positive Arens-regularity and positive symmetric Arens-regularity are equivalent. This result highlights an important difference between the theory of Banach spaces and the theory of Banach lattices. In [23], D. Leung showed that if $(e_{i})_{i}$ is the canonical basis for $\ell_{p}$, $2\leq p<\infty$, then $J(e_{i})^{\prime}$, the dual of James space $J(e_{i})$, is symmetrically Arens regular but not Arens regular. Observe that $J(e_{i})$ is not a Banach lattice since it is a non-reflexive space containing no copies of $c_{0}$ and $\ell_{1}$ (see [1, Theorem 4.71]).
Now, we prove a lemma that will be needed later in this section.
Lemma 2.2.
If $E^{*}$ has order continuous norm, then the Arens extension $\overline{A}$ of every positive symmetric $n$-linear operator $A\colon E^{n}\to F$ is also symmetric.
Proof.
Let $A\colon E\times E\to\mathbb{R}$ be a positive symmetric bilinear form. On the one hand, since $A$ is symmetric, $A=A^{t}$. On the other hand, as $E^{*}$ has order continuous norm, it follows from Proposition 2.1 that $\overline{A}=A^{t***t}$. Thus, for all $x^{**},y^{**}\in E^{**}$
$\overline{A}(x^{**},y^{**})=A^{t***t}(x^{**},y^{**})=A^{t***}(y^{**},x^{**})=A^{***}(y^{**},x^{**})=\overline{A}(y^{**},x^{**}),$
proving that $\overline{A}$ is symmetric for every positive symmetric bilinear form $A\colon E\times E\to\mathbb{R}$. Moreover, the variables in a positive symmetric $n$-linear form $A\colon E^{n}\to\mathbb{R}$ can be permuted in pairs in the extensions (see, e.g., [31, p. 52]) to get that $\overline{A}$ is symmetric.
Now, if $A\colon E^{n}\to F$ is a positive symmetric $n$-linear operator, then, for every $0\leq y^{\ast}\in F^{\ast}$, $y^{\ast}\circ A\colon E^{n}\to\mathbb{R}$ is a positive symmetric $n$-linear form, and by the above case $J_{F^{\ast}}(y^{\ast})\circ\overline{A}=\overline{y^{\ast}\circ A}$ is also a positive symmetric $n$-linear form. In particular, for every $0\leq y^{\ast\ast\ast}\in F^{\ast\ast\ast}$, we get that $0\leq y^{\ast\ast\ast}|_{F^{\ast}}\in F^{\ast}$, and hence $y^{\ast\ast\ast}\circ\overline{A}$ is symmetric. Finally, let $x_{1}^{**},\dots,x_{n}^{**}\in E^{**}$ and $\sigma\in S_{n}$. If $\overline{A}(x_{1}^{**},\dots,x_{n}^{**})\neq\overline{A}(x_{\sigma(1)}^{**},\dots,x_{\sigma(n)}^{**})$, then there exists $y^{***}\in F^{***}$ such that
$y^{***}\circ\overline{A}(x_{1}^{**},\dots,x_{n}^{**})$ $=y^{***}(\overline{A}(x_{1}^{**},\dots,x_{n}^{**}))$
$\neq y^{***}(\overline{A}(x_{\sigma(1)}^{**},\dots,x_{\sigma(n)}^{**}))=y^{***}\circ\overline{A}(x_{\sigma(1)}^{**},\dots,x_{\sigma(n)}^{**}),$
a contradiction, proving that $\overline{A}$ is symmetric.∎
The following construction is from [14, 22]. Given a bounded linear operator $\varphi\colon X^{\ast}\to Y^{\ast}$ and a symmetric $n$-linear form $A\colon X\times\cdots\times X\to\mathbb{R}$, we define $\tilde{\varphi}(A)\in\mathcal{L}(^{n}Y)$ by
$\tilde{\varphi}(A)(y_{1},\dots,y_{n})=\overline{A}(\varphi^{\ast}(J_{Y}(y_{1})),\dots,\varphi^{\ast}(J_{Y}(y_{n})))\quad\text{for all $y_{1},\dots,y_{n}\in F$}.$
In general, $\tilde{\varphi}(A)$ is not a symmetric $n$-linear form on $F$. However, for every $P\in\mathcal{P}(^{n}X)$, we can define $\overline{\varphi}(P)\in\mathcal{P}(^{n}Y)$ by
$\overline{\varphi}(P)(y)=\tilde{\varphi}(T_{P})(y,\dots,y)\quad\text{for every $y\in F$.}$
In the case $\varphi=J_{X^{\ast}}$, the morphism $\overline{\varphi}$ obtained is the Arens extension of polynomials from a Banach space $X$ to its bidual $X^{\ast\ast}$. In this case, we will use the notation $\overline{P}$ and $\overline{A}$ for $\overline{J_{X^{\ast}}}(P)$ and $\overline{J_{X^{\ast}}}(A)$, respectively. Now, for a symmetric $n$-linear map $A\colon X^{n}\to Z$, we define $\tilde{\varphi}(A)\in\mathcal{L}(^{n}Y;Z^{\ast\ast})$ by
$\tilde{\varphi}(A)(y_{1},\dots,y_{n})(z^{\ast})=\tilde{\varphi}(z^{\ast}\circ A)(y_{1},\dots,y_{n})$
for all $y_{1},\dots,y_{n}\in Y$ and $z^{\ast}\in Z^{\ast}$. Moreover, for every $P\in\mathcal{P}(^{n}X;Z)$ we may define $\overline{\varphi}(P)\in\mathcal{P}(^{n}Y;Z^{\ast\ast})$ by
$\overline{\varphi}(P)(y)=\tilde{\varphi}(T_{P})(y,\dots,y)\quad\text{for every $y\in F$}.$
The following lemma will be necessary in our discussion. Since we did not find it anywhere in the literature, we will include its proof for convenience.
Lemma 2.3.
If $A\colon E_{1}\times\cdots\times E_{n}\to F$ is a positive $n$-linear map, then its Arens extension $\overline{A}\colon E_{1}^{\ast\ast}\times\cdots\times E_{n}^{\ast\ast}\to F^{\ast\ast}$ is also positive.
Proof.
Let us recall the construction of $\overline{A}$. Let $A^{\ast}\colon F^{\ast}\times E_{1}\times\cdots\times E_{n-1}\to E_{n}^{\ast}$ be defined by
$A^{\ast}(y^{\ast},x_{1},\dots,x_{n-1})(x_{n})=y^{\ast}(A(x_{1},\dots,x_{n})).$
Then $A^{\ast\ast}\colon E_{n}^{\ast\ast}\times F^{\ast}\times E_{1}\times\cdots\times E_{n-2}\to E_{n-1}^{\ast}$ is given by
$A^{\ast\ast}(x_{n}^{\ast\ast},y^{\ast},x_{1},\dots,x_{n-2})(x_{n-1})=x_{n}^{\ast\ast}(A^{\ast}(y^{\ast},x_{1},\dots,x_{n-1}))$
Supposing that $A^{\ast(n)}\colon E_{2}^{\ast\ast}\times\cdots\times E_{n}^{\ast\ast}\times F^{\ast}\to E_{1}^{\ast}$ is defined inductively, the Arens $\overline{A}=A^{\ast(n+1)}\colon E_{1}^{\ast\ast}\times\cdots\times E_{n}^{\ast\ast}\to F^{\ast\ast}$ extension of $A$ is given by
$A^{*(n+1)}(x_{1}^{\ast\ast},\dots,x_{n}^{\ast\ast})(y^{\ast})=x_{1}^{\ast\ast}(A^{*(n)}(x_{2}^{\ast\ast},\dots,x_{n}^{\ast\ast},y^{\ast})).$
Note that for all $y^{\ast}\in(F^{\ast})^{+}$ and $x_{i}\in E_{i}^{+}$, we get that
$A^{\ast}(y^{\ast},x_{1},\dots,x_{n-1})(x_{n})=y^{\ast}(A(x_{1},\dots,x_{n}))\geq 0,$
which implies that $A^{\ast}\geq 0$. By the same reasoning, $A^{\ast\ast}\geq 0$. Inductively, we get that $\overline{A}\geq 0$.∎
Theorem 2.4.
Suppose that $E^{*}$ has order continuous norm. If $\varphi\colon E^{\ast}\to F^{\ast}$ is a lattice isomorphism, then $\overline{\varphi}\colon\mathcal{P}^{r}(^{n}E;G)\to\mathcal{P}^{r}(^{n}F;G^{\ast\ast})$ is a positive linear operator which is a linear isomorphism onto its image for every $n\in\mathbb{N}$ and every Dedekind complete Banach lattice $G$.
Proof.
We begin by proving that $\overline{\varphi}(P)\geq 0$ for every positive $n$-homogeneous polynomial $P\colon E\to G$. To see this, let $P\colon E\to G$ be a positive $n$-homogeneous polynomial. We will prove that $\tilde{\varphi(T_{P})}(y_{1},\dots,y_{n})\geq 0$ in $G^{\ast\ast}$ for all $y_{1},\dots,y_{n}\in F^{+}$ Indeed, given $y_{1},\dots,y_{n}\in F^{+}$ and $0\leq z^{\ast}\in G^{\ast}$, we have that
$\tilde{\varphi}(T_{P})(y_{1},\dots,y_{n})(z^{\ast})$ $=\tilde{\varphi}(z^{\ast}\circ T_{P})(y_{1},\dots,y_{n})$
$=\overline{z^{\ast}\circ T_{P}}(\varphi^{\ast}(J_{F}(y_{1})),\dots,\varphi^{\ast}(J_{F}(y_{n})))\geq 0$
because $\overline{z^{\ast}\circ T_{P}}\geq 0$ (due to $z^{\ast}\circ T_{P}\geq 0$), $\varphi^{\ast}\geq 0$ (due to $\varphi\geq 0$) and $J_{F}\geq 0$ (see [26, Proposition 1.4.5(ii)]), proving that $\tilde{\varphi}(T_{P})\colon F^{n}\to G^{\ast\ast}$ is a positive $n$-linear map. Consequently, $\tilde{\varphi}(T_{P})^{s}$ is also a positive $n$-linear map (see [27, Proposition 1.6]). Since $\tilde{\varphi}(T_{P})^{s}$ is the symmetric $n$-linear form associated to $\overline{\varphi}(P)$, we conclude that $\overline{\varphi}(P)$ is a positive $n$-homogeneous polynomial. Therefore, $\overline{\varphi}\colon\mathcal{P}^{r}(^{n}E;G)\to\mathcal{P}^{r}(^{n}F;G^{\ast\ast})$ is positive. Since $\varphi^{-1}\colon F^{\ast}\to E^{\ast}$ is a positive operator (see [1, Theorem 2.15]), we also have that $\overline{\varphi^{-1}}\colon\mathcal{P}^{r}(^{n}F;G^{**})\to\mathcal{P}^{r}(^{n}E;G^{****})$ is also a positive operator.
Finally, since $E^{*}$ has order continuous norm, we get by Lemma 2.2 that $\overline{T_{P}}$ is symmetric for every positive $n$-homogeneous polynomial $P\colon E\to G$. Thus, [14, Lemma 1.1(b)] yields that $\overline{\varphi^{-1}}\circ\overline{\varphi}(P)=P$ holds for every $P\in\mathcal{P}^{+}(^{n}E;G)$, and hence
$\overline{\varphi^{-1}}\circ\overline{\varphi}(P)=\overline{\varphi^{-1}}\circ\overline{\varphi}(P^{+}-P^{-})=\overline{\varphi^{-1}}\circ\overline{\varphi}(P^{+})-\overline{\varphi^{-1}}\circ\overline{\varphi}(P^{-})=P^{+}-P^{-}=P$
for every $P\in\mathcal{P}^{r}(^{n}E;G)$, which proves that $\varphi$ is a linear isomorphism onto its image. ∎
The first problem in Theorem 2.4 is that $\overline{\varphi}(P)$ may not take its values in $G$ (see [14, p. 285]). The second one is that we cannot ensure that the image of $\overline{\varphi}$ is a Banach lattice. However, when we consider polynomials taking values on a dual Banach lattice, we obtain the following.
Theorem 2.5.
If $E^{\ast}$ and $F^{\ast}$ are lattice isomorphic with at least one (and hence the other) having an order continuous norm, then $\mathcal{P}^{r}(^{n}E;G^{\ast})$ and $\mathcal{P}^{r}(^{n}F;G^{\ast})$ are lattice isomorphic for every $n\in\mathbb{N}$ and for every Banach lattice $G$.
Proof.
Let $\varphi\colon E^{\ast}\to F^{\ast}$ be the lattice isomorphism given by the assumption. We proceed as in [14]: define $\overline{\varphi_{G}}\colon\mathcal{P}(^{n}E;G^{\ast})\to\mathcal{P}(^{n}F;G^{\ast})$ by
$\overline{\varphi_{G}}(P)(y)(z)=\overline{\varphi}(J_{G}(z)\circ P)(y)$
for all $P\in\mathcal{P}(^{n}E;G^{\ast})$, $y\in F$ and $z\in G$. By the definition of $\overline{\varphi}$, we get that
$\overline{\varphi_{G}}(P)(y)(z)=\overline{P}(\varphi^{\ast}(J_{F}(y)))(z)=\overline{T_{P}}(\varphi^{\ast}(J_{F}(y)),\dots,\varphi^{\ast}(J_{F}(y)))(z).$
In particular, given $P\in\mathcal{P}^{+}(^{n}E;G^{\ast})$, the $n$-linear map $A\colon F^{n}\to G^{\ast}$ given by
$A(y_{1},\dots,y_{n})(z)=\overline{T_{P}}(\varphi^{\ast}(J_{F}(y_{1})),\dots,\varphi^{\ast}(J_{F}(y_{n})))(z)$
is positive, symmetric (because $\overline{T_{P}}$ is symmetric by the assumption and Lemma 2.2) and satisfies $\overline{\varphi_{G}}(P)(y)=A(y,\dots,y)$. Thus $\overline{\varphi_{G}}(P)\geq 0$ for every positive $n$-homogeneous polynomial $P\colon E\to G$. Therefore, we may consider the positive linear operator $\overline{\varphi_{G}}$ from $\mathcal{P}^{r}(^{n}E;G^{\ast})$ into $\mathcal{P}^{r}(^{n}F;G^{\ast})$. Analogously, since $\varphi^{-1}\colon F^{\ast}\to E^{\ast}$ is a positive operator (see [1, Theorem 2.15]), we get that $\overline{\varphi_{G}^{-1}}\colon\mathcal{P}^{r}(^{n}F;G^{\ast})\to\mathcal{P}^{r}(^{n}E;G^{\ast})$ is also a positive operator.
Finally, since $E^{*}$ and $F^{*}$ have order continuous norm, we get by the same argument used in the proof of [14, Theorem 1.3] that $\overline{\varphi_{G}^{-1}}\circ\overline{\varphi_{G}}(P)=P$ and $\overline{\varphi_{G}}\circ\overline{\varphi_{G}^{-1}}(Q)=Q$ hold for all $P\in\mathcal{P}^{+}(^{n}E;G^{\ast})$ and $Q\in\mathcal{P}^{+}(^{n}F;G^{\ast})$. Consequently,
$\overline{\varphi_{G}^{-1}}\circ\overline{\varphi_{G}}(P)=\overline{\varphi_{G}^{-1}}\circ\overline{\varphi_{G}}(P^{+}-P^{-})=\overline{\varphi_{G}^{-1}}\circ\overline{\varphi_{G}}(P^{+})-\overline{\varphi_{G}^{-1}}\circ\overline{\varphi_{G}}(P^{-})=P^{+}-P^{-}=P$
holds for every $P\in\mathcal{P}^{r}(^{n}E;G^{\ast})$, and analogously $\overline{\varphi_{G}}\circ\overline{\varphi_{G}^{-1}}(Q)=Q$ holds for every $Q\in\mathcal{P}^{r}(^{n}F;G^{\ast})$. This proves that $(\overline{\varphi}_{G})^{-1}=\overline{\varphi_{G}^{-1}}$. By [1, Theorem 2.15], we conclude that $\overline{\varphi_{G}}\colon\mathcal{P}^{r}(^{n}E;G^{\ast})\to\mathcal{P}^{r}(^{n}F;G^{\ast})$ is a lattice isomorphism. ∎
An immediate application of Theorem 2.5 is the following:
Corollary 2.6.
If $E^{*}$ and $F^{*}$ are lattice isomorphic with at least one having order continuous norm, then $\left(\widehat{\otimes}_{n,s,|\pi|}E\right)^{*}$ and $\left(\widehat{\otimes}_{n,s,|\pi|}F\right)^{*}$ are lattice isomorphic for every $n\in\mathbb{N}$.
Let us give an example:
Example 2.7.
In [21], E. Lacey constructed a nonatomic AM-space $X$ that is linearly isomorphic to $C\langle 1,\alpha\rangle$, where $\alpha\geq\omega$ is a countable ordinal and $\langle 1,\alpha\rangle$ denotes the order interval of ordinals between $1$ and $\alpha$ with the usual order topology. Moreover, there exists an isometric isomorphism $R\colon X^{*}\to\ell_{1}$ that is also order preserving, which means that $R\varphi\geq 0$ in $\ell_{1}$ if and only if $\varphi\geq 0$ in $X^{*}$. This yields that $R$ preserves the lattice structure, and hence $R$ is a lattice isomorphism (see [20, p. 5] or [19, Example 3.1]), proving that $X^{*}$ and $\ell_{1}$ are lattice isomorphic. Let us check that $c_{0}^{\ast}$ and $c^{\ast}$ are also lattice isomorphic to $\ell_{1}$:
(i) Let $S\colon\ell_{1}\to c_{0}^{*}$ be the linear isomorphism given by $S(b)(x)=\displaystyle\sum_{j=1}^{\infty}b_{j}x_{j}$ for all $b=(b_{j})_{j}\in\ell_{1}$ and $x=(x_{j})_{j}\in c_{0}$. It is easy to see that both $S$ and $S^{-1}$ are positive operators, and so $S$ is a lattice isomorphism by [1, Theorem 2.15].
(ii) Let $T\colon\ell_{1}\to c^{*}$ be the linear isomorphism given by $T(b)(x)=b_{1}\displaystyle\lim_{n\to\infty}x_{n}+\sum_{j=2}^{\infty}b_{j}x_{j-1}$ for all $b=(b_{j})_{j}\in\ell_{1}$ and $x=(x_{j})_{j}\in c$. Clearly, $T\geq 0$. The inverse map $T^{-1}\colon c^{*}\to\ell_{1}$ is such that $T^{-1}(\varphi)=(\varphi(e)-\displaystyle\sum_{j=1}^{\infty}\varphi(e_{j}),\varphi(e_{1}),\varphi(e_{2}),\dots)$, where $e=(1,1,1,\dots)$. Indeed, for every $x=(x_{j})_{j}\in c$, we have that $x=(\displaystyle\lim_{n\to\infty}x_{n})\,e\,+\,\sum_{j=1}^{\infty}(x_{j}-\lim_{n\to\infty}x_{n})e_{j},$ and hence
$T(\varphi(e)-\sum_{j=1}^{\infty}\varphi(e_{j}),\varphi(e_{1}),\varphi(e_{2}),\dots)(x)$ $=(\varphi(e)-\sum_{j=1}^{\infty}\varphi(e_{j}))\displaystyle\lim_{n\to\infty}x_{n}+\sum_{j=1}^{\infty}\varphi(e_{j})x_{j}$
$=\varphi(e)\displaystyle\lim_{n\to\infty}x_{n}+\sum_{j=1}^{\infty}(x_{j}-\lim_{n\to\infty}x_{n})\varphi(e_{j})$
$=\varphi\left(\lim_{n\to\infty}x_{n}\,e+\sum_{j=1}^{\infty}(x_{j}-\lim_{n\to\infty}x_{n})e_{j}\right)$
$=\varphi(x).$
To see that $T^{-1}\geq 0$, it is enough to observe that $\varphi(e)-\displaystyle\sum_{j=1}^{\infty}\varphi(e_{j})\geq 0$ holds for all $\varphi\geq 0$, because the positive cone is norm-closed and $\varphi(e)-\displaystyle\sum_{j=1}^{n}\varphi(e_{j})\geq 0$ for every $n\in\mathbb{N}$. Finally, since both $T$ and $T^{-1}$ are positive operators, we get from [1, Theorem 2.15] that $T$ is lattice isomorphism.
Therefore, it follows that $_{0}^{*}, cc^{*}$ and $X^{*}$ are all lattice isomorphic, while $_{0}, cc$ and $X$ are not lattice isomorphic ($c$ is not Dedekind complete, $c_{0}$ has an order continuous norm and $X$ is nonatomic, while both $c$ and $c_{0}$ are atomic). By Theorem 2.5, $\mathcal{P}^{r}(^{n}X;G^{\ast})$, $\mathcal{P}^{r}(^{n}c_{0};G^{\ast})$ and $\mathcal{P}^{r}(^{n}c;G^{\ast})$ are all lattice isomorphic for every $n\in\mathbb{N}$ and for every Banach lattice $G$.
An $n$-homogeneous polynomial $P\colon E\to F$ is said to be compact (resp. weakly-compact) if $\overline{P(B_{E})}$ is a compact (resp. weakly compact) subset of $F$. If $\mathcal{P}_{\cal K}^{+}(^{n}E;F)$ denotes the set of all positive compact $n$-homogeneous polynomials from $E$ into $F$, we define $\mathcal{P}_{\mathcal{K}}^{r}(^{n}E;F)=\text{span }\mathcal{P}_{\mathcal{K}}^{+}(^{n}E;F)$. Analogously, we define $\mathcal{P}_{\mathcal{W}}^{r}(^{n}E;F)=\text{span }\mathcal{P}_{\mathcal{W}}^{+}(^{n}E;F)$, where $\mathcal{P}_{\cal W}^{+}(^{n}E;F)$ denotes the set of all positive weakly compact $n$-homogeneous polynomials from $E$ into $F$. It was proved in [7] that under certain conditions $\mathcal{P}_{\mathcal{K}}^{r}(^{n}E;F)$ and $\mathcal{P}_{\mathcal{W}}^{r}(^{n}E;F)$ are Banach lattices with the regular norm.
Corollary 2.8.
Let $E,F,G$ be Banach lattices. Suppose that there exists a lattice isomorphism $\varphi\colon E^{\ast}\to F^{\ast}$, $E^{*}$ has order continuous norm and $G^{\ast}$ is atomic with order continuous norm or it is an AL-space. Then, $\mathcal{P}_{\mathcal{K}}^{r}(^{n}E;G^{\ast})$ and $\mathcal{P}_{\mathcal{K}}^{r}(^{n}F;G^{\ast})$ are lattice isomorphic.
Proof.
We begin recalling from [7, Example 2.3] that, under the assumption on $G^{\ast}$, $\mathcal{P}_{\mathcal{K}}^{r}(^{n}E;G^{\ast})$ and $\mathcal{P}_{\mathcal{K}}^{r}(^{n}F;G^{\ast})$ are closed sublattices of $\mathcal{P}^{r}(^{n}E;G^{\ast})$ and $\mathcal{P}^{r}(^{n}F;G^{\ast})$, respectively. Letting $\overline{\varphi_{G}}\colon\mathcal{P}^{r}(^{n}E;G^{\ast})\to\mathcal{P}^{r}(^{n}F;G^{\ast})$ be the lattice isomorphism defined in the proof of Theorem 2.5, we claim that $\overline{\varphi_{G}}(P)$ is a compact polynomial for every $P\in\mathcal{P}_{\mathcal{K}}^{+}(^{n}E;G^{\ast})$:
To prove that $\overline{\varphi_{G}}(P)$ is compact, let $(y_{k})_{k}\subset B_{F}$. In particular, $(\varphi^{\ast}(J_{F}(y_{k})))_{k}$ is a bounded sequence in $E^{\ast\ast}$. Since $P\colon E\to G^{*}$ is a compact $n$-homogeneous polynomial, it follows from [2, Proposition 2.1] that $\overline{P}\colon E^{**}\to G^{*}$ is also a compact polynomial, and hence $(\overline{P}(\varphi^{\ast}(J_{F}(y_{k}))))_{k}$ contains a convergent subsequence in $G^{\ast}$, that is $\lim_{j\to\infty}\overline{P}(\varphi^{\ast}(J_{F}(y_{k_{j}})))=g^{\ast}\in G^{*}$. Finally, since $\overline{\varphi_{G}}(P)(y_{k})=\overline{P}(\varphi^{\ast}(J_{F}(y_{k})))|_{G}$ for every $k\in\mathbb{N}$, we conclude that
$\lim_{j\to\infty}\overline{\varphi_{G}}(P)(y_{k_{j}})=\displaystyle\lim_{j\to\infty}\overline{P}(\varphi^{\ast}(J_{F}(y_{k_{j}})))|_{G}=g^{\ast\ast\ast}|_{G}\in G^{\ast},$
proving that $\overline{\varphi_{G}}(P)$ is compact. In particular, we obtain that $\overline{\varphi_{G}}\colon\mathcal{P}_{\mathcal{K}}^{r}(^{n}E;G^{\ast})\to\mathcal{P}_{\mathcal{K}}^{r}(^{n}F;G^{\ast})$ is a well defined lattice homomorphism. To conclude that $\overline{\varphi_{G}}$ is an isomorphism, just note that the same argument used above shows that $\overline{\varphi_{G}^{-1}}(Q)$ is a compact polynomial for every $Q\in\mathcal{P}_{\mathcal{K}}^{+}(^{n}F;G^{\ast})$. ∎
The following holds by applying the same argument used above replacing [7, Example 2.3] with [7, Example 2.4].
Corollary 2.9.
Let $E,F,G$ be Banach lattices. Suppose that there exists a lattice isomorphism $\varphi\colon E^{\ast}\to F^{\ast}$ and that both $E^{*}$ and $G^{\ast}$ have order continuous norms. Then, $\mathcal{P}_{\mathcal{W}}^{r}(^{n}E;G^{\ast})$ and $\mathcal{P}_{\mathcal{W}}^{r}(^{n}F;G^{\ast})$ are lattice isomorphic.
Given Banach lattices $E$, $F$. and $n\in\mathbb{N}$, we say that an $n$-homogeneous polynomial $P\in\mathcal{P}^{r}(^{n}E;F)$ is orthogonally additive if $P(x+y)=P(x)+P(y)$ whenever $|x|\wedge|y|=0$; or equivalently if $T_{P}$ is orthosymmetric, that is $T_{P}(x_{1},\dots,x_{m})=0$ whenever some pair $x_{i}$, $x_{j}$ of the arguments are disjoint, i.e. $|x_{i}|\wedge|x_{j}|=0$ (see [11, Lemma 4.1]). In the case where $F$ is Dedekind complete, the Banach lattice of all regular polynomials $P\colon E\to F$ for which $|P|$ is orthogonally additive is denoted by $\mathcal{P}_{o}^{r}(^{n}E;F)$ (see [11, Section 5]).
Theorem 2.10.
If $E^{\ast}$ and $F^{\ast}$ are lattice isomorphic with at least one (and hence the other one) having an order continuous norm, then $\mathcal{P}^{r}_{o}(^{n}E;G^{*})$ and $\mathcal{P}^{r}_{o}(^{n}F;G^{*})$ are lattice isomorphic for every $n\in\mathbb{N}$ and every Banach lattice $G$.
Proof.
Let $\varphi\colon E^{*}\to F^{*}$ be the lattice isomorphism given by the assumption. By [1, Theorem 2.20] we get that $\varphi^{*}$ is interval preserving. Since $\varphi$ is surjective, $\varphi^{*}$ is also injective. Thus, $\varphi^{*}$ is a lattice homommorphism by [30, Proposition 2.1]. Letting $\overline{\varphi_{G}}\colon\mathcal{P}^{r}(^{n}E;G^{\ast})\to\mathcal{P}^{r}(^{n}F;G^{\ast})$ be the lattice isomorphism defined in the proof of Theorem 2.5, we claim that $\overline{\varphi_{G}}(P)$ is an orthogonally additive polynomial for every $P\in\mathcal{P}_{o}^{+}(^{n}E;G^{\ast})$. Indeed, given $y_{1},\dots,y_{n}\in F$ with $|y_{i}|\wedge|y_{j}|=0$ for some $i\neq j$, we have that $|\varphi^{*}(J_{F}(y_{i}))|\wedge|\varphi^{*}(J_{F}(y_{j}))|=0$, because both $\varphi^{*}$ and $J_{F}$ preserve disjointness. Since the Arens-extension of an order bounded orthosymmetric $n$-linear map is also orthosymmetric (see [28, Corollary 5.18]), we conclude that
$T_{\overline{\varphi_{G}}(P)}(y_{1},\dots,y_{n})=\overline{T_{P}}(\varphi^{*}(J_{F}(y_{1})),\dots,\varphi^{*}(J_{F}(y_{n})))=0,$
proving that $T_{\overline{\varphi_{G}}(P)}$ is orthosymmetric. Thus, $\overline{\varphi_{G}}(P)\in\mathcal{P}_{o}^{+}(^{n}F;G^{*})$ as we claimed. Analogously, we can prove that $\overline{\varphi_{G}^{-1}}(Q)\in\mathcal{P}_{o}^{+}(^{n}E;G^{*})$ for each $Q\in\mathcal{P}_{o}^{r}(^{n}F;G^{*})$, and therefore $\overline{\varphi_{G}}\colon\mathcal{P}_{o}^{r}(^{n}E;G^{*})\to\mathcal{P}_{o}^{r}(^{n}F;G^{*})$ is a lattice isomorphism. ∎
3 Regular nuclear polynomials
Let $E$ and $F$ be two Banach lattices with $F$ being Dedekind complete. Following [4, p. 725], a bounded linear operator $T\colon E\to F$ is said to be a regular nuclear operator if it belongs to the image of the map
$\overline{\Phi}\colon E^{\ast}\,\widehat{\otimes}_{|\pi|}F\to\mathcal{L}^{r}(E;F),\quad\overline{\Phi}(x^{*}\otimes y)(x)=x^{*}(x)y.$
Fix $n\in\mathbb{N}$ and consider the positive bilinear map $A\colon\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\times F\to\mathcal{P}^{r}(^{n}E;F)$ given by $A(\theta_{n}(x^{*})\otimes y)(x)=(x^{*}(x))^{n}y$ for all $0\leq x^{*}\in E^{*}$, $y\in F^{+}$, $x\in E^{+}$ and extend by linearity and continuity. Thus, $A$ induces a positive linear operator $\overline{\Phi}_{n}\colon\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F\to\mathcal{P}^{r}(^{n}E;F)$ such that
$\overline{\Phi}_{n}(\theta_{n}(x^{*})\otimes y)(x)=(x^{*}(x))^{n}y\quad\text{for all $0\leq x^{*}\in E^{*}$, $y\in F^{+}$ and $x\in E^{+}$.}$
The elements of $\text{Im }\overline{\Phi}_{n}:=\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;F)$ are called regular nuclear $n$-homogeneous polynomials. As in [4], $\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;F)$ is a Banach lattice endowed with the norm $\|\cdot\|_{\mathcal{N},r}$ induced by the quotient norm on $\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F/\ker\overline{\Phi}$. The main objective of this section is to prove that if $G$ is a Dedekind complete Banach lattice with the lattice approximation property and both $E^{*}$ and $F^{*}$ are lattice isomorphic, then $\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;G)$ and $\mathcal{P}_{\mathcal{N}}^{r}(^{n}F;G)$ are lattice isomorphic. To do this, we recall from [4] that a principal pair of a Banach lattice $E$ is an ordered pair $(G,g)$, in which, $G$ is a finite dimensional vector sublattice of $E$ and $g$ is a strong order unit of $G$, i.e. for each $y\in G$, exists $\lambda>0$ with $|y|\leq\lambda g$. A Banach lattice $E$ is said to have the lattice approximation property (LAP in short) if for every $\varepsilon>0$ and every sequence $(E_{n},e_{n})$ of principal pairs of $E$ such that $\|e_{n}\|\to 0$ there exists finite rank operator $T\colon E\to E$ such that $\||(T-id)_{E_{n}}|(e_{n})\|<\varepsilon$ for every $n\in\mathbb{N}$.
The following lemma will be necessary to achieve our objective.
Lemma 3.1.
Let $E$ and $F$ be two Banach lattices with $F$ being Dedekind complete. For every $w\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$, there exist a sequence $(F_{k})_{k}$ of finite-dimensional vector sublattices of $F$ and a sequence $(w_{k})_{k}$ in $\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$ such that $\|w\|_{|\pi|}=\sum_{k=1}^{\infty}\|w_{k}\|_{|\pi|}<\infty$, $w=\sum_{k=1}^{\infty}w_{k}$, $w_{k}\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\otimes F_{k}$ and $\|w_{k}\|_{\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F_{k}}\leq 2\|w_{k}\|_{|\pi|}+2^{1-k}$ for every $k\in\mathbb{N}$. Moreover, $\|w_{k}\|_{\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F_{k}}\leq\frac{6}{2^{k}}$ holds for every $k\in\mathbb{N}$ whenever $\|w\|_{|\pi|}=1$.
Proof.
Let $w\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$ be given. By [4, Lemma 3.5], there exist a finite dimensional vector sublattice $F_{1}$ of $F$ and a vector $w_{1}\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\otimes F_{1}$ such that $\|w-w_{1}\|_{|\pi|}\leq 2^{-1}$ and $\|w_{1}\|_{\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F_{1}}\leq 2\|w_{1}\|_{|\pi|}+1$. Using [4, Lemma 3.5] again, there exist a finite dimensional vector sublattice $F_{2}$ of $F$ and a vector $w_{2}\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\otimes F_{k}$ such that $\|(w-w_{1})-w_{2}\|_{|\pi|}\leq 2^{-2}$ and $\|w_{2}\|_{\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F_{2}}\leq 2\|w_{2}\|_{|\pi|}+2^{-1}$. Proceeding inductively, we may construct a sequence $(F_{k})_{k}$ of finite dimensional vectors sublattices of $F$ and a sequence $(w_{k})_{k}\subset\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$ such that $w_{k}\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\otimes F_{k}$, $\|(w-w_{1}-\cdots-w_{k-1})-w_{k}\|_{|\pi|}\leq 2^{-k}$ and $\|w_{k}\|_{\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F_{k}}\leq 2\|w_{k}\|_{|\pi|}+2^{1-k}$ for every $k>1$. Thus,
$\left|\|w\|_{|\pi|}-\sum_{i=1}^{k}\|w_{i}\|_{|\pi|}\right|\leq\|w-\sum_{i=1}^{k}w_{i}\|_{|\pi|}\leq 2^{-k}$
holds for every $k\in\mathbb{N}$. By letting $k\to\infty$, we get that $\|w\|_{|\pi|}=\sum_{i=1}^{\infty}\|w_{i}\|_{|\pi|}<\infty$ and $w=\displaystyle\sum_{k=1}^{\infty}w_{k}$. We also notice that $\|w_{k+1}\|_{|\pi|}\leq\|w\|_{|\pi|}-\sum_{i=1}^{k}\|w_{i}\|_{|\pi|}\leq 2^{-k}$ holds for every $k\in\mathbb{N}$, and so
$\|w_{k}\|_{\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F_{k}}$ $\leq 2\|w_{k}\|_{|\pi|}+2^{1-k}\leq 2\,2^{-(k-1)}+2^{1-k}$
$=2\,2^{-k+1}+2^{1-k}=2^{1-k}(2+1)=\frac{6}{2^{k}}$
for every $k\geq 2.$ Assuming $\|w\|_{|\pi|}=1$, for $k=1$, we have that
$\|w_{1}\|_{\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F_{1}}\leq 2\|w_{1}\|_{|\pi|}+1\leq 2\|w\|_{|\pi|}+1\leq 3=\displaystyle\frac{6}{2^{1}}$
and we are done. ∎
It was proved in [4] that, whenever $F$ is a Dedekind complete Banach lattice with the LAP, then $\overline{\Phi}\colon E^{\ast}\,\widehat{\otimes}_{|\pi|}F\to\mathcal{N}^{r}(E;F)$ is one-to-one, and hence an isometry, for every Banach lattice $E$. The same argument used in Blanco’s proof can be adapt to obtain a polynomial version of this result that will be presented in our next theorem. First, we recall that $\mathcal{P}_{f}(^{n}E;F)$ denotes the vector space of all finite rank $n$-homogeneous polynomials from $E$ into $F$.
Theorem 3.2.
If $F$ is a Dedekind complete Banach lattice with the LAP, then the map $\overline{\Phi}_{n}\colon\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F\to\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;F)$ is injective, and hence an isometry.
Proof.
Suppose for the sake of contradiction that $\overline{\Phi}_{n}$ is not injective. Thus, there exists $w\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$ with $\|w\|_{|\pi|}=1$ such that $\overline{\Phi}_{n}(w)=0$. Let $(F_{k})_{k}$ and $(w_{k})_{k}$ be as in Lemma 3.1. Since, for every $k\in\mathbb{N}$, $w_{k}\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\otimes F_{k}$ and $\|w_{k}\|_{\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F_{k}}\leq 2\|w_{k}\|_{|\pi|}+2^{1-k}\leq\displaystyle\frac{6}{2^{k}}$, there are a positive elements $z_{k,1},\dots,z_{k,m_{k}}\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}$ and positive vectors $y_{k,1},\dots,y_{k,m_{k}}\in F_{k}$ such that
$|w_{k}|\leq\displaystyle\sum_{j=1}^{m_{k}}z_{k,j}\otimes y_{k,j}\quad\text{ and}\quad\displaystyle\sum_{j=1}^{m_{k}}\|z_{k,j}\|_{s,|\pi|}\|y_{k,j}\|\leq\frac{6}{2^{k}}.$
For each $k\in\mathbb{N}$ and $j=1,\dots,m_{k}$, we let
$u_{k,j}=k\|y_{k,j}\|z_{k,j}\quad\text{and}\quad v_{k,j}\displaystyle=\left\{\begin{matrix}\frac{y_{k,j}}{k\|y_{k,j}\|},\,\text{if }y_{k,j}\neq 0;\\ 0,\,\text{if }y_{k,j}=0.\end{matrix}\right$.
Thus, $0\leq u_{k,1},\dots,u_{k,m_{k}}$ $\in$ $\widehat{\otimes}_{n,s,|\pi|}E^{\ast}$ and $0\leq y_{k,1},\dots,y_{k,m_{k}}\in F_{k}$ satisfy
$|w_{k}|\leq\sum_{j=1}^{m_{k}}z_{k,j}\otimes y_{k,j}=\sum_{j=1}^{m_{k}}u_{k,j}\otimes v_{k,j},\quad\lim_{k\to\infty}\max_{1\leq j\leq m_{k}}\|v_{k,j}\|=\lim_{k\to\infty}\frac{1}{k}=0,$
and
$\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\|u_{k,j}\|_{s,|\pi|}=\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}k\|z_{k,j}\|_{s,|\pi|}\|y_{k,j}\|\leq\sum_{k=1}^{\infty}\frac{6k}{2^{k}}<\infty.$
Defining $F_{k,j}=\left\{{y\in F_{k}}\,:\,{|y|\leq m\,v_{k,j}\,\text{ for some }m\in\mathbb{N}}\right\}$ for all $k,j\in\mathbb{N}$, we get that $((F_{k,j})_{k,j},v_{k,j})$ is a sequence of principal pairs of $F$.
Let $\varepsilon>0$. As $F$ has the LAP, there exists a finite rank operator $T\colon F\to F$ such that $\||(T-id)|_{F_{k,j}}|(v_{k,j})\|<\varepsilon$ for all $k,j\in\mathbb{N}$. Defining $\Psi\colon\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F\to\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$ by $\Psi(z\otimes y)=z\otimes T(y)$ for all $z\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}$ and $y\in F$, we get that $\overline{\Phi}_{n}\circ\Psi(u)=T\circ(\overline{\Phi}_{n}(u))$ for every $u\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$, and hence $\overline{\Phi}_{n}(\Psi(w))=T\circ(\overline{\Phi}_{n}(w))=0$. As $\Psi(w_{k})\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\otimes T(F)$ for every $k\in\mathbb{N}$, $\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\otimes T(F)$ is closed in $\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$ and $\Psi(w)=\displaystyle\sum_{k=1}^{\infty}\Psi(w_{k})$, we get that $\Psi(w)\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\otimes T(F)$, which yields that $0=\overline{\Phi}_{n}(\Psi(w))=\Phi_{n}(\Psi(w))$, where $\Phi_{n}$ is the following isomorphism of vector spaces:
$\Phi_{n}\colon\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\otimes F\to\mathcal{P}_{f}(^{n}E;F),\quad\Phi_{n}(\theta_{n}(x^{*})\otimes y)=(x^{*}(x))^{n}y.$
Now, we may proceed as in [4, p.730] to conclude that $\|w\|_{|\pi|}\leq\varepsilon\sum_{k=1}^{\infty}\sum_{j=1}^{n_{k}}\|u_{k,j}\|_{s,|\pi|}$, and hence $w=0$, a contradiction. ∎
Corollary 3.3.
Let $E,F,G$ be Banach lattices such that $G$ is Dedekind complete with the LAP. If $E^{*}$ and $F^{*}$ are lattice isomorphic, then $\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;G^{\ast})$ and $\mathcal{P}_{\mathcal{N}}^{r}(^{n}F;G^{\ast})$ are lattice isomorphic.
Proof.
If $E^{*}$ and $F^{*}$ are lattice isomorphic, then $\widehat{\otimes}_{k,s,|\pi|}E^{\ast}$ and $\widehat{\otimes}_{k,s,|\pi|}F^{\ast}$ are lattice isomorphic. Indeed, if $\varphi\colon E^{*}\to F^{*}$ is a lattice isomorphism, then $P(x^{*})=\theta_{n}(\varphi(x^{*}))$ defines a positive $n$-homogeneous polynomial from $E$ into $\widehat{\otimes}_{k,s,|\pi|}F^{\ast}$, and so $P^{\otimes}\colon\widehat{\otimes}_{k,s,|\pi|}E^{\ast}\to\widehat{\otimes}_{k,s,|\pi|}F^{\ast}$ is a positive linear operator such that $P^{\otimes}(\theta_{n}(x^{*}))=\theta_{n}(\varphi(x^{*}))$ for every $x^{*}\in E^{*}$. Analogously, there exists a positive operator $S\colon\widehat{\otimes}_{k,s,|\pi|}F^{\ast}\to\widehat{\otimes}_{k,s,|\pi|}E^{\ast}$ such that $S(y^{*})=\theta_{n}(\varphi^{-1}(y^{*}))$ for every $y^{*}\in F^{*}$. Finally, since
$S\circ P^{\otimes}(\theta_{n}(x^{*}))=S(\theta_{n}(\varphi(x^{*})))=\theta_{n}(\varphi^{-1}(\varphi(x^{*})))=\theta_{n}(x^{*})$
holds for every $x^{*}\in E^{*}$, we obtain that $S\circ P^{\otimes}(z)=z$ for every $z\in\widehat{\otimes}_{k,s,|\pi|}E^{\ast}$. The same idea also proves that $P^{\otimes}\circ S(z)=z$ for every $z\in\widehat{\otimes}_{k,s,|\pi|}F^{\ast}$. Thus, $(P^{\otimes})^{-1}=S$, and by [1, Theorem 2.15] $P^{\otimes}$ is a lattice isomorphism. An adaptation of the above argument can be used to prove that $\widehat{\otimes}_{k,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}G$ is lattice isomorphic to $\widehat{\otimes}_{k,s,|\pi|}F^{\ast}\,\widehat{\otimes}_{|\pi|}G$ . Now, by applying Theorem 3.2, we get that $\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;G)$ and $\mathcal{P}_{\mathcal{N}}^{r}(^{n}F;G)$ are lattice isomorphic. ∎
C. Boyd
Department of Mathematics and Statistics
University College Dublin
Belfield – Dublin 4 – Ireland
e-mail: christopher.boyd@ucd.ie
V. C. C. Miranda
Centro de Matemática, Computação e Cognição
Universidade Federal do ABC
09.210-580 – Santo André – Brazil.
e-mail: colferaiv@gmail.com