### v2

# Lattice isomorphic Banach lattices of polynomials

Abstract
We study Díaz-Dineen’s problem for regular homogeneous vector-valued polynomials. In particular, we prove that if $E^{*}$ and $F^{*}$ are lattice isomorphic with at least one having order continuous norm, then $\mathcal{P}^{r}(^{n}E;G^{*})$ and $\mathcal{P}^{r}(^{n}F;G^{*})$ are lattice isomorphic for every $n\in\mathbb{N}$ and every Banach lattice $G$. We also study the analogous problem for the classes of regular compact, regular weakly compact, orthogonally additive and regular nuclear polynomials.
1 Introduction and background
In [18], J. C. Díaz and S. Dineen posed the following question: If $X$ and $Y$ are dual isomorphic Banach spaces, that is $X^{\ast}$ and $Y^{\ast}$ are isomorphic as Banach spaces, are $\mathcal{P}(^{n}X)$ and $\mathcal{P}(^{n}Y)$ isomorphic as Banach spaces? Although it is not known if the above question has a positive answer, some partial answers were presented in [16, 18, 29]. In particular, J. C. Díaz and S. Dineen proved that the problem has a positive solution whenever $X^{\ast}$ has both the Schur property and the approximation property (see [18, Proposition 4]). In [16], F. Cabello Sánchez, J. Castillo and R. García generalized Díaz-Dineen results by providing a positive solution for this problem whenever $X$ is Arens-regular or whenever $X$ and $Y$ are stable. The same result was proven, independently, by S. Lassalle and I. Zalduendo in [29] who also addressed the same question concerning other classes of polynomials. Further, in [17], D. Carando and S. Lassalle extended these results to spaces of vector-valued homogeneous polynomials.
Motivated by the increasing interest in studying homogeneous polynomials on Banach lattices (see, e.g., [6, 7, 8, 9, 10, 11, 13, 24, 31]), the following question arises naturally: Given two dual isomorphic Banach lattices $E$ and $F$, that is $E^{\ast}$ and $F^{\ast}$ are lattice isomorphic, under which conditions are $\mathcal{P}^{r}(^{n}E)$ and $\mathcal{P}^{r}(^{n}F)$ lattice isomorphic for every $n\in\mathbb{N}$? By adapting the proofs from [17] and [29], we were able to provide a positive answer whenever $E^{*}$ and $F^{*}$ have order continuous norms. In particular, we prove that if $E^{\ast}$ and $F^{\ast}$ are lattice isomorphic with at least one (and hence the other) having order continuous norm, then $\mathcal{P}^{r}(^{n}E;G^{*})$ and $\mathcal{P}^{r}(^{n}F;G^{*})$ are lattice isomorphic for every $n\in\mathbb{N}$ and every Banach lattice $G$. We observe that, in our setting, the assumption of Arens regularity is replaced with the requirement that the norm on the dual is order continuous. This replacement is justified by the fact that a positive version of Arens regularity for $E$ is equivalent to the order continuity of the norm on $E^{*}$. Furthermore, as we will demonstrate in Proposition 2.1, positive Arens regularity coincides with classical Arens regularity and with positive symmetric Arens regularity, highlighting a noteworthy contrast with the Banach space case.
As an application, we obtain conditions so that $\mathcal{P}_{\cal K}^{r}(^{n}E;G^{*})$ (resp. $\mathcal{P}_{\cal W}^{r}(^{n}E;G^{*})$) and $\mathcal{P}_{\cal K}^{r}(^{n}F;G^{*})$ (resp. $\mathcal{P}_{\cal W}^{r}(^{n}F;G^{*})$) are isomorphic, where $\mathcal{P}_{\cal K}^{r}$ and $\mathcal{P}_{\cal W}^{r}$ denote, respectively, the classes of regular compact and regular weakly compact polynomials. We also consider the class of orthogonally additive polynomials. In Section 3, we investigate this problem for a new class of polynomials. Specifically, inspired by ideas from [4], we define the class of regular nuclear $n$-homogeneous polynomials, denoted by $\mathcal{P}_{\cal N}^{r}$, and prove that if $G$ is Dedekind complete and has the lattice approximation property (LAP), and $E^{*}$ and $F^{*}$ are lattice isomorphic both having the LAP, then $\mathcal{P}_{\cal N}^{r}(^{n}E;G)$ and $\mathcal{P}_{\cal N}^{r}(^{n}F;G)$ are lattice isomorphic for every $n\in\mathbb{N}$.
Before proceeding, we introduce some definitions and notation. Throughout this paper, $X$ and $Y$ denote Banach spaces, $E$ and $F$ denote Banach lattices, $E^{+}$, $B_{E}$ and $S_{E}$ denote, respectively, the positive cone, the closed unit ball, and the unit sphere of $E$. Given Banach spaces $X_{1},\dots,X_{n}$ and $Y$, the Banach space of all continuous $n$-linear operators $A\colon X_{1}\times\cdots\times X_{n}\to Y$ is denoted by $\mathcal{L}(X_{1},\dots,X_{n};Y)$. In the case that $F=\mathbb{K}$, we simply write $\mathcal{L}(X_{1},\dots,X_{n})$. Given Banach lattices $E_{1},\dots,E_{n}$ and $F$, a $n$-linear operator $A\colon E_{1}\times\cdots\times E_{n}\to F$ is said to be positive if $A(x_{1},\dots,x_{n})\geq 0$ for all $x_{1}\in E_{1}^{+},\dots,x_{n}\in E_{n}^{+}$. The difference of two positive $n$-linear operators is called a regular $n$-linear operator, and the set of all regular $n$-linear operators from $E_{1}\times\cdots\times E_{n}$ into $F$ is denoted by $\mathcal{L}^{r}(E_{1},\dots,E_{n};F)$. Whenever $F$ is Dedekind complete, $\mathcal{L}^{r}(E_{1},\dots,E_{n};F)$ is a Banach lattice with the regular norm $\|A\|_{r}=\||A|\|$, where $|A|$ denotes the absolute value of the regular $n$-linear operator $A\colon E_{1}\times\cdots\times E_{n}\to F$. The positive projective tensor product, or the Fremlin tensor product of $E_{1},\dots,E_{n}$, denoted by $E_{1}\widehat{\otimes}_{|\pi|}\cdots\widehat{\otimes}_{|\pi|}E_{n}$, is the Banach lattice generated by considering the completion of the vector lattice tensor product $E_{1}\overline{\otimes}\cdots\overline{\otimes}\,E_{n}$ with respect the positive projective tensor norm:
$\|u\|_{|\pi|}=\inf\left\{{\sum_{i=1}^{k}\|x_{i}^{1}\|\cdots\|x_{i}^{n}\|}\,:\,{|u|\leq\sum_{i=1}^{n}x_{i}^{1}\otimes\cdots\otimes x_{i}^{n},\,x_{i}^{j}\in E_{j}^{+}}\right\}.$
For each $T\in\mathcal{L}^{r}(E_{1},\dots,E_{m};F)$, there exists a unique bounded linear operator $T^{\otimes}$ from $E_{1}\widehat{\otimes}_{|\pi|}\cdots\widehat{\otimes}_{|\pi|}E_{n}$ into $F$, called the linearization of $T$, such that $T^{\otimes}(x_{1}\otimes\cdots\otimes x_{n})=T(x_{1},\dots,x_{n})$ for all $x_{1}\in E_{1},\dots,x_{n}\in E_{n}$. Moreover, the map
$T\in\mathcal{L}^{r}(E_{1},\dots,E_{m};F)\mapsto T^{\otimes}\in\mathcal{L}^{r}(E_{1}\widehat{\otimes}_{|\pi|}\cdots\widehat{\otimes}_{|\pi|}E_{n};F)$
is an isometric isomorphism and a lattice hommomorphism (see [14, Proposition 3.3]).
An $n$-homogeneous polynomial between Riesz spaces $P\colon E\to F$ is positive if its associated symmetric multilinear operator $T_{P}\colon E^{n}\to F$ is positive. The difference of two positive $n$-homogeneous polynomials is called a regular homogeneous polynomial, and the set of such polynomials is denoted by $\mathcal{P}^{r}(^{n}E,F)$. When $F$ is the scalar field we simply write $\mathcal{P}^{r}(^{n}E)$. If $E$ and $F$ are Banach lattices with $F$ Dedekind complete, then $\mathcal{P}^{r}(^{n}E,F)$ is a Banach lattice with the regular norm $\|P\|_{r}=\||P|\|$, where $|P|$ denotes the absolute value of the regular $n$-homogeneous polynomial $P\colon E\to F$. The $n$-fold positive projective symmetric tensor product of $E$, denoted by $\widehat{\otimes}_{n,s,|\pi|}E$, is a Banach lattice endowed with the positive projective symmetric tensor norm:
$\|u\|_{s,|\pi|}=\inf\left\{{\sum_{i=1}^{k}\|x_{i}\|^{n}}\,:\,{|u|\leq\sum_{i=1}^{k}\otimes^{n}x_{i},\,x_{i}\in E^{+}}\right\},$
where $\otimes^{n}x=x\,\otimes\stackrel{{\scriptstyle n}}{{\cdots}}\otimes\,x$ for every $x\in E$. Let $\theta_{n}\colon E\to\widehat{\otimes}_{n,s,|\pi|}E$ be the canonical $n$-homogeneous polynomial given by $\theta_{n}(x)=\otimes^{n}x$. We note that $\theta_{n}$ is a lattice homomorphism. For every $P\in\mathcal{P}^{r}(^{n}E,F)$ there exists a unique regular linear operator $P^{\otimes}\colon\widehat{\otimes}_{n,s,|\pi|}E\to F$, called the linearization of $P$, such that $P(x)=P^{\otimes}(\theta_{n}(x))$ for every $x\in E$. The operator
$\Phi\colon\mathcal{L}^{r}\left(\widehat{\otimes}_{n,s,|\pi|}E;F\right)\to\mathcal{P}^{r}(^{n}E;F)~,~\Phi(T)=T\circ\theta_{n},\quad T\in\mathcal{L}^{r}(\widehat{\otimes}_{n,s,|\pi|}E;F),$
is an isometric isomorphism and a lattice homomorphism [14, Proposition 3.4].
We refer the reader to [1, 33] for background on Banach lattices, to [21] for Banach space theory, to [19] for polynomials on Banach spaces, to [14, 32] for positive/regular multilinear operators or polynomials, and to [14, 22, 23] for tensor products between vector lattices.
2 Díaz-Dineen problem for regular polynomials
We begin this section recalling that the first Arens adjoint of $A\in\mathcal{L}(X_{1},\cdots,X_{n};Y)$ is the $n$-linear map $A^{*}\colon F^{*}\times E_{1}\times\cdots\times E_{n-1}\to E_{n}^{*}$ defined by
$A^{*}(y^{*},x_{1},\dots,x_{n-1})(x_{n})=y^{*}(A(x_{1},\dots,x_{n})).$
The second Arens adjoint $A^{**}$ of $A$ is the first Arens adjoint of $A^{*}$, that is the $n$-linear map $A^{**}\colon E_{n}^{**}\times F^{*}\times E_{1}\times\cdots\times E_{n-2}\to E_{n-1}^{*}$ given by
$A^{**}(x_{n}^{**},y^{*},x_{1},\dots,x_{n-1})(x_{n}^{*})=x_{n}^{**}(A^{*}(y^{*},x_{1},\dots,x_{n-1})).$
This process is repeated until we obtain $\overline{A}:=A^{(n+1)*}\colon E_{1}^{**}\times\cdots\times E_{n}^{**}\to F^{**}$, the $(n+1)$-Arens adjoint of $A$ which is called the Arens-extension of $A$ (see [9, Section 2] for more details). For the remainder of the paper, we shall call this extension the Arens extension, bearing in mind that it coincides with its Aron-Berner extension (see [3]). Whenever $X=Y$, the transpose of $A$ is the bilinear map $A^{t}\colon X\times X\to Z$ given by $A^{t}(x,y)=A(y,x)$ for all $x,y\in X$. Still under the assumption that $X=Y$, AA is said to be Arens-regular (or simply regular) if $A^{t***t}=A^{***}$. It is important to note that we have two different concepts of regularity for multilinear operators, one in the lattice sense presented in the Introduction and the one in the extension sense. It follows from [37, Theorem 2.2] that a bilinear form $A\colon X\times X\to\mathbb{R}$ is Arens-regular if and only if the bounded linear operator $S_{A}\colon X\to X^{*}$ given by $S_{A}(x)(y)=A(x,y)$ is weakly compact.
Recall that a Banach space $X$ is called Arens-regular if all linear operators from $X$ into $X^{*}$ are weakly compact. To our knowledge, the first appearance of a positive version of Arens-regularity was in [15], when G. Buskes and R. Page gave necessary and sufficient conditions so that every positive bilinear form $E\times E\to\mathbb{R}$ is Arens-regular. The following result collects the characterizations obtained so far by G. Buskes, R. Page, G. Emmanuele, E. Saab and P. Saab, and proves that positive Arens-regularity does not give rise to a genuinely new concept as it is equivalent to classical Arens-regularity.
Proposition 2.1.
For a Banach lattice $E$, the following are equivalent:
(1) $E^{\ast}$ has order continuous norm.
(2) $E$ does not contain a sublattice isomorphic to $\ell_{1}$.
(3) Every positive bilinear map $A\colon E\times E\to\mathbb{R}$ is Arens-regular.
(4) Every positive operator $T\colon E\to E^{*}$ is weakly compact.
(5) Every positive symmetric bilinear map $A\colon E\times E\to\mathbb{R}$ is Arens-regular.
(6) $E$ is Arens-regular.
Proof.
[…]
$A\in\mathcal{L}^{+}(E,E;\mathbb{R})\mapsto S_{A}\in\mathcal{L}^{+}(E;E^{\ast}).$
[…]
$B(x,y)=\sum_{j\text{ even}}\sum_{k=1}^{j-1}x_{j}y_{k}\,+\sum_{k\text{ even}}\sum_{j=1}^{k-1}x_{j}y_{k},\quad x=(x_{i})_{i},y=(y_{i})_{i}\in\ell_{1}.$
Since $B$ is positive, we may consider the positive symmetric bilinear map $A\colon E\times E\to\mathbb{R}$ given by $A(x,y)=B(\Pi(x),\Pi(y))$ for all $x,y\in E$. Let us see that $A$ cannot be Arens-regular, showing that its associated operator $S_{A}\colon E\to E^{*}$ is not weakly compact: as $\{S_{B}(e_{2n}):n\in\mathbb{N}\}$ is not a weakly precompact subset of $\ell_{1}^{*}=\ell_{\infty}$ (see [3, p. 83]), there exists $\varphi\in\ell_{\infty}$ such that $(\varphi(S_{B}(e_{2n})))_{n}$ does not have any Cauchy subsequence. If $x_{n}\in E$ is such that $\Pi(x_{n})=e_{2n}$, then $(\varphi(S_{B}(\Pi(x_{n}))))_{n}$ does not have any Cauchy subsequence. By extending $\varphi\colon\ell_{1}\to\mathbb{R}$ to $E$ in such a way that it vanishes on the complement, we obtain that $\{\Pi^{*}\circ S_{B}\circ\Pi(x_{n}):n\in\mathbb{N}\}$ is not weakly precompact in $E^{*}$, and since $S_{A}=\Pi^{*}\circ S_{B}\circ\Pi$, we conclude that $S_{A}$ is not weakly compact.
[…]
It is worth mentioning that Proposition 2.1 shows that the notions of Arens-regularity, positive Arens-regularity, and positive symmetric Arens-regularity are all equivalent in the Banach lattice setting. This result highlights an important difference between the theory of Banach spaces and the theory of Banach lattices. In [30], D. Leung showed that if $(e_{i})_{i}$ is the canonical basis for $\ell_{p}$, $2\leq p<\infty$, then $J(e_{i})^{\prime}$, the dual of James space $J(e_{i})$, is symmetrically Arens regular but not Arens regular. Observe that $J(e_{i})$ is not a Banach lattice since it is a non-reflexive space containing no copies of $c_{0}$ or $\ell_{1}$ (see [1, Theorem 4.71]).
Our next result is an immediate consequence of Proposition 2.1 and [39, p. 52].
Corollary 2.2.
If $E^{*}$ has order continuous norm, then the Arens extension $\overline{A}$ of every positive symmetric $n$-linear operator $A\colon E^{n}\to F$ is also symmetric.
The following construction is from [17, 29]. Given a bounded linear operator $\varphi\colon X^{\ast}\to Y^{\ast}$ and a symmetric $n$-linear form $A\colon X\times\cdots\times X\to\mathbb{R}$, we define $\tilde{\varphi}(A)\in\mathcal{L}(^{n}Y)$ by
$\tilde{\varphi}(A)(y_{1},\dots,y_{n})=\overline{A}(\varphi^{\ast}(J_{Y}(y_{1})),\dots,\varphi^{\ast}(J_{Y}(y_{n})))\quad\text{for all $y_{1},\dots,y_{n}\in F$}.$
In general, $\tilde{\varphi}(A)$ is not a symmetric $n$-linear form on $F$. However, for every $P\in\mathcal{P}(^{n}X)$, we can define $\overline{\varphi}(P)\in\mathcal{P}(^{n}Y)$ by
$\overline{\varphi}(P)(y)=\tilde{\varphi}(T_{P})(y,\dots,y)\quad\text{for every $y\in F$.}$
In the case $\varphi=J_{X^{\ast}}$, the morphism $\overline{\varphi}$ obtained is the Arens extension of polynomials from a Banach space $X$ to its bidual $X^{\ast\ast}$. In this case, we will use the notation $\overline{P}$ and $\overline{A}$ for $\overline{J_{X^{\ast}}}(P)$ and $\overline{J_{X^{\ast}}}(A)$, respectively. Now, for a symmetric $n$-linear map $A\colon X^{n}\to Z$, we define $\tilde{\varphi}(A)\in\mathcal{L}(^{n}Y;Z^{\ast\ast})$ by
$\tilde{\varphi}(A)(y_{1},\dots,y_{n})(z^{\ast})=\tilde{\varphi}(z^{\ast}\circ A)(y_{1},\dots,y_{n})$
for all $y_{1},\dots,y_{n}\in Y$ and $z^{\ast}\in Z^{\ast}$. Moreover, for every $P\in\mathcal{P}(^{n}X;Z)$ we may define $\overline{\varphi}(P)\in\mathcal{P}(^{n}Y;Z^{\ast\ast})$ by
$\overline{\varphi}(P)(y)=\tilde{\varphi}(T_{P})(y,\dots,y)\quad\text{for every $y\in F$}.$
In the proof of our next results, we will use the fact that the Arens extension of every positive $n$-linear map is also positive [5, Theorem 2.2(c)]. A proof of this result can be found in the doctoral dissertation of L. Garcia [26, Theorem 2.1.3].
Theorem 2.3.
Suppose that $E^{*}$ has order continuous norm. If $\varphi\colon E^{\ast}\to F^{\ast}$ is a lattice isomorphism, then $\overline{\varphi}\colon\mathcal{P}^{r}(^{n}E;G)\to\mathcal{P}^{r}(^{n}F;G^{\ast\ast})$ is a positive linear operator which is a linear isomorphism onto its image for every $n\in\mathbb{N}$ and every Dedekind complete Banach lattice $G$.
Proof.
We begin by proving that $\overline{\varphi}(P)\geq 0$ for every positive $n$-homogeneous polynomial $P\colon E\to G$. To see this, let $P\colon E\to G$ be a positive $n$-homogeneous polynomial. We will prove that $\tilde{\varphi}(T_{P})(y_{1},\dots,y_{n})\geq 0$ in $G^{\ast\ast}$ for all $y_{1},\dots,y_{n}\in F^{+}$. Indeed, given $y_{1},\dots,y_{n}\in F^{+}$ and $0\leq z^{\ast}\in G^{\ast}$, we have that
$\tilde{\varphi}(T_{P})(y_{1},\dots,y_{n})(z^{\ast})$ $=\tilde{\varphi}(z^{\ast}\circ T_{P})(y_{1},\dots,y_{n})$
$=\overline{z^{\ast}\circ T_{P}}(\varphi^{\ast}(J_{F}(y_{1})),\dots,\varphi^{\ast}(J_{F}(y_{n})))\geq 0$
because $\overline{z^{\ast}\circ T_{P}}\geq 0$ (due to $z^{\ast}\circ T_{P}\geq 0$), $\varphi^{\ast}\geq 0$ (due to $\varphi\geq 0$) and $J_{F}\geq 0$ (see [33, Proposition 1.4.5(ii)]). This proves that $\tilde{\varphi}(T_{P})\colon F^{n}\to G^{\ast\ast}$ is a positive $n$-linear map, and consequently $\tilde{\varphi}(T_{P})^{s}$ is also a positive $n$-linear map (see [34, Proposition 1.6]). Since $\tilde{\varphi}(T_{P})^{s}$ is the symmetric $n$-linear form associated to $\overline{\varphi}(P)$, we conclude that $\overline{\varphi}(P)$ is a positive $n$-homogeneous polynomial. Therefore, $\overline{\varphi}\colon\mathcal{P}^{r}(^{n}E;G)\to\mathcal{P}^{r}(^{n}F;G^{\ast\ast})$ is positive. Since $\varphi^{-1}\colon F^{\ast}\to E^{\ast}$ is a positive operator (see [1, Theorem 2.15]), we also have that $\overline{\varphi^{-1}}\colon\mathcal{P}^{r}(^{n}F;G^{**})\to\mathcal{P}^{r}(^{n}E;G^{****})$ is also a positive operator.
Finally, since $E^{*}$ has order continuous norm, we get by Corollary 2.2 that $\overline{T_{P}}$ is symmetric for every positive $n$-homogeneous polynomial $P\colon E\to G$. Thus, [17, Lemma 1.1(b)] yields that $\overline{\varphi^{-1}}\circ\overline{\varphi}(P)=P$ holds for every $P\in\mathcal{P}^{+}(^{n}E;G)$, and hence
$\overline{\varphi^{-1}}\circ\overline{\varphi}(P)=\overline{\varphi^{-1}}\circ\overline{\varphi}(P^{+}-P^{-})=\overline{\varphi^{-1}}\circ\overline{\varphi}(P^{+})-\overline{\varphi^{-1}}\circ\overline{\varphi}(P^{-})=P^{+}-P^{-}=P$
for every $P\in\mathcal{P}^{r}(^{n}E;G)$, which proves that $\varphi$ is a linear isomorphism onto its image. ∎
The first limitation of Theorem 2.3 is that $\overline{\varphi}(P)$ may not take its values in $G$ (see [17, p. 285]). The second is that we cannot ensure that the image of $\overline{\varphi}$ is a Banach lattice. However, when we consider polynomials taking values on a dual Banach lattice, we obtain the following.
Theorem 2.4.
If $E^{\ast}$ and $F^{\ast}$ are lattice isomorphic with at least one (and hence the other) having an order continuous norm, then $\mathcal{P}^{r}(^{n}E;G^{\ast})$ and $\mathcal{P}^{r}(^{n}F;G^{\ast})$ are lattice isomorphic for every $n\in\mathbb{N}$ and for every Banach lattice $G$.
Proof.
Let $\varphi\colon E^{\ast}\to F^{\ast}$ be the lattice isomorphism given by the assumption. We proceed as in [17]: define $\overline{\varphi_{G}}\colon\mathcal{P}(^{n}E;G^{\ast})\to\mathcal{P}(^{n}F;G^{\ast})$ by
$\overline{\varphi_{G}}(P)(y)(z)=\overline{\varphi}(J_{G}(z)\circ P)(y)$
for all $P\in\mathcal{P}(^{n}E;G^{\ast})$, $y\in F$ and $z\in G$. By the definition of $\overline{\varphi}$, we get that
$\overline{\varphi_{G}}(P)(y)(z)=\overline{P}(\varphi^{\ast}(J_{F}(y)))(z)=\overline{T_{P}}(\varphi^{\ast}(J_{F}(y)),\dots,\varphi^{\ast}(J_{F}(y)))(z).$
In particular, given $P\in\mathcal{P}^{+}(^{n}E;G^{\ast})$, the $n$-linear map $A\colon F^{n}\to G^{\ast}$ given by
$A(y_{1},\dots,y_{n})(z)=\overline{T_{P}}(\varphi^{\ast}(J_{F}(y_{1})),\dots,\varphi^{\ast}(J_{F}(y_{n})))(z)$
is positive, symmetric (because $\overline{T_{P}}$ is symmetric by the assumption and Corollary 2.2) and satisfies $\overline{\varphi_{G}}(P)(y)=A(y,\dots,y)$. Thus $\overline{\varphi_{G}}(P)\geq 0$ for every positive $n$-homogeneous polynomial $P\colon E\to G$. Therefore, we may consider the positive linear operator $\overline{\varphi_{G}}$ from $\mathcal{P}^{r}(^{n}E;G^{\ast})$ into $\mathcal{P}^{r}(^{n}F;G^{\ast})$. Analogously, since $\varphi^{-1}\colon F^{\ast}\to E^{\ast}$ is a positive operator (see [1, Theorem 2.15]), we get that $\overline{\varphi_{G}^{-1}}\colon\mathcal{P}^{r}(^{n}F;G^{\ast})\to\mathcal{P}^{r}(^{n}E;G^{\ast})$ is also a positive operator.
Finally, since $E^{*}$ and $F^{*}$ have order continuous norm, we get by the same argument used in the proof of [17, Theorem 1.3] that $\overline{\varphi_{G}^{-1}}\circ\overline{\varphi_{G}}(P)=P$ and $\overline{\varphi_{G}}\circ\overline{\varphi_{G}^{-1}}(Q)=Q$ hold for all $P\in\mathcal{P}^{+}(^{n}E;G^{\ast})$ and $Q\in\mathcal{P}^{+}(^{n}F;G^{\ast})$. Consequently,
$\overline{\varphi_{G}^{-1}}\circ\overline{\varphi_{G}}(P)=\overline{\varphi_{G}^{-1}}\circ\overline{\varphi_{G}}(P^{+}-P^{-})=\overline{\varphi_{G}^{-1}}\circ\overline{\varphi_{G}}(P^{+})-\overline{\varphi_{G}^{-1}}\circ\overline{\varphi_{G}}(P^{-})=P^{+}-P^{-}=P$
holds for every $P\in\mathcal{P}^{r}(^{n}E;G^{\ast})$, and analogously $\overline{\varphi_{G}}\circ\overline{\varphi_{G}^{-1}}(Q)=Q$ holds for every $Q\in\mathcal{P}^{r}(^{n}F;G^{\ast})$. This proves that $(\overline{\varphi}_{G})^{-1}=\overline{\varphi_{G}^{-1}}$. By [1, Theorem 2.15], we conclude that $\overline{\varphi_{G}}\colon\mathcal{P}^{r}(^{n}E;G^{\ast})\to\mathcal{P}^{r}(^{n}F;G^{\ast})$ is a lattice isomorphism. ∎
An immediate application of Theorem 2.4 is the following:
Corollary 2.5.
If $E^{*}$ and $F^{*}$ are lattice isomorphic with at least one having order continuous norm, then $\left(\widehat{\otimes}_{n,s,|\pi|}E\right)^{*}$ and $\left(\widehat{\otimes}_{n,s,|\pi|}F\right)^{*}$ are lattice isomorphic for every $n\in\mathbb{N}$.
Let us give an example:
Example 2.6.
In [28], E. Lacey constructed a nonatomic AM-space $X$ that is linearly isomorphic to $C\langle 1,\alpha\rangle$, where $\alpha\geq\omega$ is a countable ordinal and $\langle 1,\alpha\rangle$ denotes the order interval of ordinals between $1$ and $\alpha$ with the usual order topology. Moreover, there exists an isometric isomorphism $R\colon X^{*}\to\ell_{1}$ that is also order preserving, which means that $R\varphi\geq 0$ in $\ell_{1}$ if and only if $\varphi\geq 0$ in $X^{*}$. This means that $R$ preserves the lattice structure, and hence $R$ is a lattice isomorphism (see [27, p. 5] or [25, Example 3.1]), proving that $X^{*}$ and $\ell_{1}$ are lattice isomorphic. Let us check that $c_{0}^{\ast}$ and $c^{\ast}$ are also lattice isomorphic to $\ell_{1}$:
(i) Let $S\colon\ell_{1}\to c_{0}^{*}$ be the linear isomorphism given by $S(b)(x)=\displaystyle\sum_{j=1}^{\infty}b_{j}x_{j}$ for all $b=(b_{j})_{j}\in\ell_{1}$ and $x=(x_{j})_{j}\in c_{0}$. It is easy to see that both $S$ and $S^{-1}$ are positive operators, and so $S$ is a lattice isomorphism by [1, Theorem 2.15].
(ii) Let $T\colon\ell_{1}\to c^{*}$ be the linear isomorphism given by $T(b)(x)=b_{1}\displaystyle\lim_{n\to\infty}x_{n}+\sum_{j=2}^{\infty}b_{j}x_{j-1}$ for all $b=(b_{j})_{j}\in\ell_{1}$ and $x=(x_{j})_{j}\in c$. Clearly, $T\geq 0$. The inverse map $T^{-1}\colon c^{*}\to\ell_{1}$ is given by $T^{-1}(\varphi)=(\varphi(e)-\displaystyle\sum_{j=1}^{\infty}\varphi(e_{j}),\varphi(e_{1}),\varphi(e_{2}),\dots)$, where $e=(1,1,1,\dots)$. Indeed, for every $x=(x_{j})_{j}\in c$, we have that $x=(\displaystyle\lim_{n\to\infty}x_{n})\,e\,+\,\sum_{j=1}^{\infty}(x_{j}-\lim_{n\to\infty}x_{n})e_{j},$ and hence
$T(\varphi(e)-\sum_{j=1}^{\infty}\varphi(e_{j}),\varphi(e_{1}),\varphi(e_{2}),\dots)(x)$ $=(\varphi(e)-\sum_{j=1}^{\infty}\varphi(e_{j}))\displaystyle\lim_{n\to\infty}x_{n}+\sum_{j=1}^{\infty}\varphi(e_{j})x_{j}$
$=\varphi(e)\displaystyle\lim_{n\to\infty}x_{n}+\sum_{j=1}^{\infty}(x_{j}-\lim_{n\to\infty}x_{n})\varphi(e_{j})$
$=\varphi\left(\lim_{n\to\infty}x_{n}\,e+\sum_{j=1}^{\infty}(x_{j}-\lim_{n\to\infty}x_{n})e_{j}\right)$
$=\varphi(x).$
To see that $T^{-1}\geq 0$, it is enough to observe that $\varphi(e)-\displaystyle\sum_{j=1}^{\infty}\varphi(e_{j})\geq 0$ holds for all $\varphi\geq 0$, because the positive cone is norm-closed and $\varphi(e)-\displaystyle\sum_{j=1}^{n}\varphi(e_{j})\geq 0$ for every $n\in\mathbb{N}$. Finally, since both $T$ and $T^{-1}$ are positive operators, we get from [1, Theorem 2.15] that $T$ is lattice isomorphism.
Therefore, it follows that $_{0}^{*}, cc^{*}$ and $X^{*}$ are all lattice isomorphic, while $_{0}, cc$ and $X$ are not lattice isomorphic ($c$ is not Dedekind complete, $c_{0}$ has an order continuous norm and $X$ is nonatomic, while both $c$ and $c_{0}$ are atomic). By Theorem 2.4, $\mathcal{P}^{r}(^{n}X;G^{\ast})$, $\mathcal{P}^{r}(^{n}c_{0};G^{\ast})$ and $\mathcal{P}^{r}(^{n}c;G^{\ast})$ are all lattice isomorphic for every $n\in\mathbb{N}$ and for every Banach lattice $G$.
An $n$-homogeneous polynomial $P\colon E\to F$ is said to be compact (resp. weakly-compact) if $\overline{P(B_{E})}$ is a compact (resp. weakly compact) subset of $F$. If $\mathcal{P}_{\cal K}^{+}(^{n}E;F)$ denotes the set of all positive compact $n$-homogeneous polynomials from $E$ into $F$, we define $\mathcal{P}_{\mathcal{K}}^{r}(^{n}E;F)=\text{span }\mathcal{P}_{\mathcal{K}}^{+}(^{n}E;F)$. Analogously, we define $\mathcal{P}_{\mathcal{W}}^{r}(^{n}E;F)=\text{span }\mathcal{P}_{\mathcal{W}}^{+}(^{n}E;F)$, where $\mathcal{P}_{\cal W}^{+}(^{n}E;F)$ denotes the set of all positive weakly compact $n$-homogeneous polynomials from $E$ into $F$. It was proved in [8] that under certain conditions $\mathcal{P}_{\mathcal{K}}^{r}(^{n}E;F)$ and $\mathcal{P}_{\mathcal{W}}^{r}(^{n}E;F)$ are Banach lattices with the regular norm.
Proposition 2.7.
Let $E,F,G$ be Banach lattices. Suppose that there exists a lattice isomorphism $\varphi\colon E^{\ast}\to F^{\ast}$, $E^{*}$ has order continuous norm and $G^{\ast}$ is atomic with order continuous norm or it is an AL-space. Then, $\mathcal{P}_{\mathcal{K}}^{r}(^{n}E;G^{\ast})$ and $\mathcal{P}_{\mathcal{K}}^{r}(^{n}F;G^{\ast})$ are lattice isomorphic.
Proof.
We begin recalling from [8, Example 2.3] that, under the assumption on $G^{\ast}$, $\mathcal{P}_{\mathcal{K}}^{r}(^{n}E;G^{\ast})$ and $\mathcal{P}_{\mathcal{K}}^{r}(^{n}F;G^{\ast})$ are closed sublattices of $\mathcal{P}^{r}(^{n}E;G^{\ast})$ and $\mathcal{P}^{r}(^{n}F;G^{\ast})$, respectively. Letting $\overline{\varphi_{G}}\colon\mathcal{P}^{r}(^{n}E;G^{\ast})\to\mathcal{P}^{r}(^{n}F;G^{\ast})$ be the lattice isomorphism defined in the proof of Theorem 2.4, we claim that $\overline{\varphi_{G}}(P)$ is a compact polynomial for every $P\in\mathcal{P}_{\mathcal{K}}^{+}(^{n}E;G^{\ast})$:
To prove that $\overline{\varphi_{G}}(P)$ is compact, let $(y_{k})_{k}\subset B_{F}$. In particular, $(\varphi^{\ast}(J_{F}(y_{k})))_{k}$ is a bounded sequence in $E^{\ast\ast}$. Since $P\colon E\to G^{*}$ is a compact $n$-homogeneous polynomial, it follows from [2, Proposition 2.1] that $\overline{P}\colon E^{**}\to G^{*}$ is also a compact polynomial, and hence $(\overline{P}(\varphi^{\ast}(J_{F}(y_{k}))))_{k}$ contains a convergent subsequence in $G^{\ast}$, that is $\lim_{j\to\infty}\overline{P}(\varphi^{\ast}(J_{F}(y_{k_{j}})))=g^{\ast}\in G^{*}$. Finally, since $\overline{\varphi_{G}}(P)(y_{k})=\overline{P}(\varphi^{\ast}(J_{F}(y_{k})))|_{G}$ for every $k\in\mathbb{N}$, we conclude that
$\lim_{j\to\infty}\overline{\varphi_{G}}(P)(y_{k_{j}})=\displaystyle\lim_{j\to\infty}\overline{P}(\varphi^{\ast}(J_{F}(y_{k_{j}})))|_{G}=g^{\ast}\in G^{\ast},$
proving that $\overline{\varphi_{G}}(P)$ is compact. In particular, we obtain that $\overline{\varphi_{G}}\colon\mathcal{P}_{\mathcal{K}}^{r}(^{n}E;G^{\ast})\to\mathcal{P}_{\mathcal{K}}^{r}(^{n}F;G^{\ast})$ is a well defined lattice homomorphism. To conclude that $\overline{\varphi_{G}}$ is an isomorphism, just note that the same argument used above shows that $\overline{\varphi_{G}^{-1}}(Q)$ is a compact polynomial for every $Q\in\mathcal{P}_{\mathcal{K}}^{+}(^{n}F;G^{\ast})$. ∎
The following holds by applying the same argument used above replacing [8, Example 2.3] with [8, Example 2.4].
Proposition 2.8.
Let $E,F,G$ be Banach lattices. Suppose that there exists a lattice isomorphism $\varphi\colon E^{\ast}\to F^{\ast}$ and that both $E^{*}$ and $G^{\ast}$ have order continuous norms. Then, $\mathcal{P}_{\mathcal{W}}^{r}(^{n}E;G^{\ast})$ and $\mathcal{P}_{\mathcal{W}}^{r}(^{n}F;G^{\ast})$ are lattice isomorphic.
Given Banach lattices $E$, $F$, and $n\in\mathbb{N}$, we say that an $n$-homogeneous polynomial $P\in\mathcal{P}^{r}(^{n}E;F)$ is orthogonally additive if $P(x+y)=P(x)+P(y)$ whenever $|x|\wedge|y|=0$; or equivalently if $T_{P}$ is orthosymmetric, that is $T_{P}(x_{1},\dots,x_{m})=0$ whenever some pair $x_{i}$, $x_{j}$ of the arguments are disjoint, i.e. $|x_{i}|\wedge|x_{j}|=0$ (see [14, Lemma 4.1]). In the case where $F$ is Dedekind complete, the Banach lattice of all regular polynomials $P\colon E\to F$ for which $|P|$ is orthogonally additive is denoted by $\mathcal{P}_{o}^{r}(^{n}E;F)$ (see [14, Section 5]).
Theorem 2.9.
If $E^{\ast}$ and $F^{\ast}$ are lattice isomorphic with at least one (and hence the other one) having an order continuous norm, then $\mathcal{P}^{r}_{o}(^{n}E;G^{*})$ and $\mathcal{P}^{r}_{o}(^{n}F;G^{*})$ are lattice isomorphic for every $n\in\mathbb{N}$ and every Banach lattice $G$.
Proof.
Let $\varphi\colon E^{*}\to F^{*}$ be the lattice isomorphism given by the assumption. By [1, Theorem 2.20] we get that $\varphi^{*}$ is interval preserving. Since $\varphi$ is surjective, $\varphi^{*}$ is also injective. Thus, $\varphi^{*}$ is a lattice homomorphism by [38, Proposition 2.1]. Letting $\overline{\varphi_{G}}\colon\mathcal{P}^{r}(^{n}E;G^{\ast})\to\mathcal{P}^{r}(^{n}F;G^{\ast})$ be the lattice isomorphism defined in the proof of Theorem 2.4, we claim that $\overline{\varphi_{G}}(P)$ is an orthogonally additive polynomial for every $P\in\mathcal{P}_{o}^{+}(^{n}E;G^{\ast})$. Indeed, given $y_{1},\dots,y_{n}\in F$ with $|y_{i}|\wedge|y_{j}|=0$ for some $i\neq j$, we have that $|\varphi^{*}(J_{F}(y_{i}))|\wedge|\varphi^{*}(J_{F}(y_{j}))|=0$, because both $\varphi^{*}$ and $J_{F}$ preserve disjointness. Since the Arens-extension of an order bounded orthosymmetric $n$-linear map is also orthosymmetric (see [35, Corollary 5.18]), we conclude that
$T_{\overline{\varphi_{G}}(P)}(y_{1},\dots,y_{n})=\overline{T_{P}}(\varphi^{*}(J_{F}(y_{1})),\dots,\varphi^{*}(J_{F}(y_{n})))=0,$
proving that $T_{\overline{\varphi_{G}}(P)}$ is orthosymmetric. Thus, $\overline{\varphi_{G}}(P)\in\mathcal{P}_{o}^{+}(^{n}F;G^{*})$ as we claimed. Analogously, we can prove that $\overline{\varphi_{G}^{-1}}(Q)\in\mathcal{P}_{o}^{+}(^{n}E;G^{*})$ for each $Q\in\mathcal{P}_{o}^{+}(^{n}F;G^{*})$, and therefore $\overline{\varphi_{G}}\colon\mathcal{P}_{o}^{r}(^{n}E;G^{*})\to\mathcal{P}_{o}^{r}(^{n}F;G^{*})$ is a lattice isomorphism. ∎
3 Regular nuclear polynomials
Let $E$ and $F$ be two Banach lattices with $F$ being Dedekind complete. Following [4, p. 725], a bounded linear operator $T\colon E\to F$ is said to be a regular nuclear operator if it belongs to the image of the map
$\overline{\Phi}\colon E^{\ast}\,\widehat{\otimes}_{|\pi|}F\to\mathcal{L}^{r}(E;F),\quad\overline{\Phi}(x^{*}\otimes y)(x)=x^{*}(x)y.$
To present a polynomial version of such class, fix $n\in\mathbb{N}$ and let $\phi_{1},\dots,\phi_{n}\in(E^{*})^{+}$ be given. If $A:E\times\overset{n}{\cdots}\times E\to\mathbb{R}$ is the $n$-linear operator defined by
$A(x_{1},\dots,x_{n})=\phi_{1}(x_{1})\cdots\phi_{n}(x_{n}),$
its symmetrization is given by
$A^{s}(x_{1},\dots,x_{n})=\frac{1}{n!}\sum_{\sigma\in S_{n}}A(x_{\sigma(1)},\dots,x_{\sigma(n)})=\frac{1}{n!}\sum_{\sigma\in S_{n}}\phi_{1}(x_{\sigma(1)})\cdots\phi_{n}(x_{\sigma(n)}).$
Since $\phi_{1},\dots,\phi_{n}\geq 0$, we get that $A^{s}\geq 0$. Thus $P_{\phi_{1},\dots,\phi_{n}}(x):=A(x,\dots,x)=\phi_{1}(x)\cdots\phi_{n}(x)$ defines a positive $n$-homogeneous polynomial on $E$, and so we can define a positive symmetric $n$-linear map
$(\phi_{1},\dots,\phi_{n})\in E^{*}\times\cdots\times E^{*}\mapsto P_{\phi_{1},\dots,\phi_{n}}\in\mathcal{P}^{r}(^{n}E).$
Its range is contained in $\mathcal{P}^{r}(^{n}E)$. For instance, taking $n=2$,
$\phi_{1}\cdot\phi_{2}=(\phi_{1}^{+}-\phi_{1}^{-})(\phi_{2}^{+}-\phi_{2}^{-})=(\phi_{1}^{+}\phi_{2}^{+}+\phi_{1}^{-}\phi_{2}^{-})-(\phi_{1}^{+}\phi_{2}^{-}+\phi_{1}^{-}\phi_{2}^{+}).$
Now, we can consider the positive bilinear map $B\colon\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\times F\to\mathcal{P}^{r}(^{n}E;F)$ given by $B(\theta_{n}(x^{*})\otimes y)(x)=T(\theta_{n}(x^{*}))y=(x^{*}(x))^{n}y$ for all $0\leq x^{*}\in E^{*}$, $y\in F^{+}$, $x\in E^{+}$. Hence, $B$ induces a positive linear operator $\overline{\Phi}_{n}\colon\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F\to\mathcal{P}^{r}(^{n}E;F)$ such that
$\overline{\Phi}_{n}(\theta_{n}(x^{*})\otimes y)(x)=(x^{*}(x))^{n}y\quad\text{for all $0\leq x^{*}\in E^{*}$, $y\in F^{+}$ and $x\in E^{+}$.}$
The elements of $\text{Im }\overline{\Phi}_{n}:=\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;F)$ are called regular nuclear $n$-homogeneous polynomials. As in [4], $\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;F)$ is a Banach lattice endowed with the norm $\|\cdot\|_{\mathcal{N},r}$ induced by the quotient norm on $\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F/\ker\overline{\Phi}$. The main objective of this section is to prove that if $G$ is a Dedekind complete Banach lattice with the lattice approximation property, and both $E^{*}$ and $F^{*}$ have the lattice approximation property and are lattice isomorphic, then $\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;G)$ and $\mathcal{P}_{\mathcal{N}}^{r}(^{n}F;G)$ are lattice isomorphic. To do this, we recall from [4] that a principal pair of a Banach lattice $E$ is an ordered pair $(G,g)$, in which, $G$ is a finite dimensional vector sublattice of $E$ and $g$ is a strong order unit of $G$, i.e. for each $y\in G$, exists $\lambda>0$ with $|y|\leq\lambda g$. A Banach lattice $E$ is said to have the lattice approximation property (LAP in short) if for every $\varepsilon>0$ and every sequence $(E_{n},e_{n})$ of principal pairs of $E$ such that $\|e_{n}\|\to 0$ there exists finite rank operator $T\colon E\to E$ such that $\||(T-id)_{E_{n}}|(e_{n})\|<\varepsilon$ for every $n\in\mathbb{N}$.
The following lemma, which can be seen as a generalization of [4, Lemma 3.5], will be necessary to achieve our objective.
Lemma 3.1.
Let $E$ and $F$ be Dedekind complete Banach lattices and $w\in\widehat{\otimes}_{n,s,|\pi|}E\,\widehat{\otimes}_{|\pi|}\,F$. For each $0<\varepsilon\leq 1$, there exist finite-dimensional vector sublattices $E_{0}\subset E$ and $F_{0}\subset F$ and a vector $\widetilde{w}\in\otimes_{n,s}E_{0}\,\otimes F_{0}$ such that $\|w-\tilde{w}\|_{s,|\pi|}\leq\varepsilon$ and $\|\tilde{w}\|_{\otimes_{n,s}E_{0}\widehat{\otimes}_{|\pi|}F_{0}}\leq(1+\varepsilon)^{n}\|\tilde{w}\|_{s,|\pi|}+2^{n}\varepsilon$
Proof.
Since $\widehat{\otimes}_{n,s,|\pi|}E\otimes\,F$ is dense in $\widehat{\otimes}_{n,s,|\pi|}E\,\widehat{\otimes}_{|\pi|}\,F$ there exists $w_{0}=\displaystyle\sum_{j=1}^{m}u_{j}\otimes y_{j}\in\widehat{\otimes}_{n,s,|\pi|}E\otimes\,F$ such that $\|w-w_{0}\|_{|\pi|}\leq\dfrac{\varepsilon}{4}$. Without loss of generality, we assume that $u_{j}\neq 0$ and $y_{j}\neq 0$ for every $j=1,\dots,n$. Now, from the density of ${\otimes}_{n,s}E$ in $\widehat{\otimes}_{n,s,|\pi|}E$, for each $j=1,\dots,m$, there exists $u_{j,0}:=\displaystyle\sum_{i=1}^{m_{j}}\lambda_{i,j}\theta_{n}(x_{i}^{j})$ such that $\|u_{j}-u_{j,0}\|_{s,|\pi|}\leq\dfrac{\varepsilon}{4m\max\{\|y_{1}\|,\dots,\|y_{m}\|\}}$. Thus $u:=\sum_{j=1}^{m}\sum_{i=1}^{m_{j}}\lambda_{i,j}\theta_{n}(x_{i}^{j})\otimes y_{j}\in{\otimes}_{n,s}E\otimes F$ satisfies
$\|w-u\|_{|\pi|}$ $\leq\|w-w_{0}\|_{|\pi|}+\|w_{0}-u\|_{|\pi|}$
$\leq\frac{\varepsilon}{4}+\|\sum_{j=1}^{m}(u_{j}\otimes y_{j}-\sum_{i=1}^{m_{j}}\lambda_{i,j}\theta_{n}(x_{i}^{j})\otimes y_{j})\|_{|\pi|}$
$=\frac{\varepsilon}{4}+\|\sum_{j=1}^{m}(u_{j}-\sum_{i=1}^{m_{j}}\ \lambda_{i,j}\theta_{n}(x_{i}^{j}))\otimes y_{j}\|_{|\pi|}$
$\leq\frac{\varepsilon}{4}+\sum_{j=1}^{m}\|u_{j}-\sum_{i=1}^{m_{j}}\lambda_{i,j}\theta_{n}(x_{i}^{j})\|_{s,|\pi|}\|y_{j}\|\leq\frac{\varepsilon}{2}.$
Also, by the definition of the positive projective tensor norm, there exist $z_{1},\dots,z_{p}\in(\widehat{\otimes}_{n,s,|\pi|}E)^{+}$ and $v_{1},\dots,v_{p}\in F^{+}$ such that
$|u|\leq\sum_{i=1}^{p}z_{i}\otimes v_{i}\,\text{ and }\,\sum_{i=1}^{p}\|z_{i}\|_{s,|\pi|}\|v_{i}\|\leq\|u\|_{|\pi|}+\frac{\varepsilon}{4}.$
Without loss of generality, we assume that $v_{i}\neq 0$ for every $i=1,\dots,p$. For each $i=1,\dots,p$, we apply the definition of the positive projective symmetric tensor norm to obtain $z_{i}^{1},\dots,z_{i}^{p_{i}}\in E^{+}$ such that
$z_{i}=|z_{i}|\leq\sum_{k=1}^{p_{i}}\theta_{n}(z_{i}^{k})\,\text{ and }\,\sum_{k=1}^{p_{i}}\|z_{i}^{k}\|^{n}\leq\|z_{i}\|_{s,|\pi|}+\frac{\varepsilon}{4p\max\{\|v_{1}\|,\dots,\|v_{p}\|\}}.$
Therefore,
$|u|\leq\sum_{i=1}^{p}z_{i}\otimes v_{i}\leq\sum_{i=1}^{p}(\sum_{k=1}^{p_{i}}\theta_{n}(z_{i}^{k}))\otimes v_{i}=\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}\theta_{n}(z_{i}^{k})\otimes v_{i}$
and
$\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}\|z_{i}^{k}\|^{n}\|v_{i}\|$ $=\sum_{i=1}^{p}\left(\sum_{k=1}^{p_{i}}\|z_{i}^{k}\|^{n}\right)\|v_{i}\|\leq\sum_{i=1}^{p}\left(\|z_{i}\|_{s,|\pi|}+\frac{\varepsilon}{4p\max\{\|v_{1}\|,\dots,\|v_{p}\|\}}\right)\|v_{i}\|$
$=\sum_{k=1}^{p}\|z_{i}\|_{s,|\pi|}\|v_{i}\|+\frac{\varepsilon}{4}\leq\|u\|_{|\pi|}+\frac{\varepsilon}{4}+\frac{\varepsilon}{4}=\|u\|_{|\pi|}+\frac{\varepsilon}{2}.$
Let
$E_{0}:={\rm span}\{x_{1}^{j},\dots,x_{m_{1}}^{j},\,z_{i}^{1},\dots,z_{i}^{p_{i}}\,:j=1,\dots,m,\,i=1,\dots,p\}\subset E$
and
$F_{0}:={\rm span}\{y_{1},\dots,y_{m},\,v_{1},\dots,v_{p}\}\subset F.$
Since $E$ and $F$ are Dedekind complete, by [4, Proposition 2.1], for each $\delta>0$, there are finite-dimensional vector sublattices $E_{\delta}\subset E$, $F_{\delta}\subset F$, and linear maps
$U_{\delta}:E_{0}\to E_{\delta},\,V_{\delta}:F_{0}\to F_{\delta}$
such that
$\|U_{\delta}(x)-x\|\leq\delta\|x\|\,\text{for every $x\in E_{0}$}\,\text{ and }\,{\|V_{\delta}(y)-y\|\leq\delta\|y\|}\,\text{for every $y\in F_{0}$}.$
For each $\delta>0$, set $v^{\delta}=\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}\theta_{n}(U_{\delta}(z_{i}^{k}))\otimes V_{\delta}(v_{i})$. We claim that $v^{\delta}\longrightarrow v:=\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}\theta_{n}(z_{i}^{k})\otimes v_{i}$ as $\delta\longrightarrow 0$. We have
$\|v-v^{\delta}\|_{|\pi|}$ $\leq\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}\|\theta_{n}(U_{\delta}(z_{i}^{k}))\otimes V_{\delta}(v_{i})-\theta_{n}(z_{i}^{k})\otimes v_{i}\|_{|\pi|}$
$\leq\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}[\|\theta_{n}(U_{\delta}(z_{i}^{k}))\|_{s,|\pi|}\|V_{\delta}(v_{i})-v_{i}\|+\|\theta_{n}(U_{\delta}(z_{i}^{k}))-\theta_{n}(z_{i}^{k})\|_{s,|\pi|}\|v_{i}\|]$
$\leq\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}[\|U_{\delta}(z_{i}^{k})\|^{n}\|V_{\delta}(v_{i})-v_{i}\|+n\,\max\{\|U_{\delta}(z_{i}^{k})\|,\|z_{i}^{k}\|\}^{n-1}\|U_{\delta}(z_{i}^{k})-z_{i}^{k}\|\|v_{i}\|]$
$\leq\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}[M^{n}\delta\|v_{i}\|+nM^{n}\delta\|z_{i}^{k}\|]$
$\leq\delta\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}[M^{n+1}+nM^{n+1}],$
provided $\delta\leq 1$, where $M=\displaystyle\max_{i=1,\dots,p}\{\|U_{\delta}(z_{i}^{1})\|,\dots,\|U_{\delta}(z_{i}^{p_{i}})\|,\|z_{i}^{1}\|,\dots,\|z_{i}^{p_{i}}\|,\|v_{i}\|\}$. Letting $\delta\to 0$, we obtain that $\|v-v^{\delta}\|_{|\pi|}\longrightarrow 0$. Analogously, we have that
$u^{\delta}:=\sum_{j=1}^{m}\sum_{i=1}^{m_{j}}\lambda_{i,j}\theta_{n}(U_{\delta}(x_{i}^{j}))\otimes V_{\delta}(y_{j})\longrightarrow u.$
The continuity of the lattice operations yield that
$w^{\delta}:=v^{\delta}\wedge((-v^{\delta})\vee u^{\delta})\to v\wedge((-v)\vee u)=u\quad\text{as }\delta\to 0.$
Now, we choose $0<\delta_{0}<\min\{1,\varepsilon\}$ such that $\|u-w^{\delta_{0}}\|_{|\pi|}\leq\frac{\varepsilon}{2}$. We will prove that $E_{\delta_{0}}$, $F_{\delta_{0}}$ and $\widetilde{w}:=w^{\delta_{0}}$ satisfy this lemma hypothesis. First, we note that since $E_{\delta_{0}}$ and $F_{\delta_{0}}$ are finite-dimensional vector lattices, $\otimes_{n,s}E_{\delta_{0}}\otimes F_{\delta_{0}}$ is finite-dimensional vector lattice, and so $\tilde{w}\in\otimes_{n,s}E_{\delta_{0}}\otimes F_{\delta_{0}}$. Secondly, we have that
$\|w-\tilde{w}\|_{|\pi|}\leq\|w-u\|_{|\pi|}+\|u-\tilde{w}\|_{|\pi|}\leq\varepsilon.$
Moreover, since
$|\widetilde{w}|\leq|v^{\delta_{0}}|\leq\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}|\theta_{n}(U_{\delta_{0}}(z_{i}^{k}))|\otimes|V_{\delta_{0}}(v_{i})|=\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}\theta_{n}(|U_{\delta_{0}}(z_{i}^{k})|)\otimes|V_{\delta_{0}}(v_{i})|,$
we get
$\|\widetilde{w}\|_{\otimes_{n,s}E_{\delta_{0}}\widehat{\otimes}_{|\pi|}F_{\delta_{0}}}$ $\leq\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}\|\theta_{n}(U_{\delta_{0}}(z_{i}^{k}))\|_{s,|\pi|}\|V_{\delta_{0}}(v_{i})\|$
$=\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}\|U_{\delta_{0}}(z_{i}^{k})\|^{n}\|V_{\delta_{0}}(v_{i})\|\leq\sum_{i=1}^{p}\sum_{k=1}^{p_{i}}(1+\delta_{0})^{n+1}\|z_{i}^{k}\|^{n}\|v_{i}\|$
$\leq(1+\delta_{0})^{n+1}(\|u\|_{|\pi|}+\frac{\varepsilon}{2})\leq(1+\varepsilon)^{n+1}(\|\widetilde{w}\|_{|\pi|}+\varepsilon)\leq(1+\varepsilon)^{n+1}\|\widetilde{w}\|_{|\pi|}+2^{n}\varepsilon,$
and we are done. ∎
We will also need the following technical lemma.
Lemma 3.2.
Let $E$ and $F$ be two Dedekind complete Banach lattices. For every $w\in\widehat{\otimes}_{n,s,|\pi|}E\,\widehat{\otimes}_{|\pi|}F$, there exist sequences $(E_{k})_{k}$ and $(F_{k})_{k}$ of finite-dimensional vector sublattices of $E$ and $F$, respectively, and a sequence $(w_{k})_{k}$ in $\widehat{\otimes}_{n,s,|\pi|}E\,\widehat{\otimes}_{|\pi|}F$ such that $\|w\|_{|\pi|}=\sum_{k=1}^{\infty}\|w_{k}\|_{|\pi|}<\infty$, $w=\sum_{k=1}^{\infty}w_{k}$, $w_{k}\in\otimes_{n,s}E_{k}\otimes F_{k}$ and $\|w_{k}\|_{\otimes_{n,s}E_{k}\,\widehat{\otimes}_{|\pi|}F_{k}}\leq 2^{n}\|w_{k}\|_{|\pi|}+2^{n-k}$ for every $k\in\mathbb{N}$. Moreover, $\|w_{k}\|_{\otimes_{n,s}E_{k}\,\widehat{\otimes}_{|\pi|}F_{k}}\leq\frac{2^{n}\cdot 3}{2^{k}}$ holds for every $k\in\mathbb{N}$ whenever $\|w\|_{|\pi|}=1$.
Proof.
Let $w\in\widehat{\otimes}_{n,s,|\pi|}E\,\widehat{\otimes}_{|\pi|}F$ be given. By Lemma 3.1, there exist finite dimensional vector sublattices $E_{1}\subset E$ and $F_{1}\subset F$, and a vector $w_{1}\in\otimes_{n,s}E_{1}\,\otimes F_{1}$ such that $\|w-w_{1}\|_{|\pi|}\leq 2^{-1}$ and
$\|w_{1}\|_{\otimes_{n,s}E_{1}\,\widehat{\otimes}_{|\pi|}F_{1}}\leq\left(1+\frac{1}{2}\right)^{n}\|w_{1}\|_{|\pi|}+2^{n}\dfrac{1}{2}\leq 2^{n}\|w_{1}\|_{|\pi|}+2^{n-1}.$
Using Lemma 3.1 again, there exist finite dimensional vector sublattices $E_{2}\subset E$ and $F_{2}\subset F$, and a vector $w_{2}\in\otimes_{n,s}E_{2}\,\otimes F_{2}$ such that $\|(w-w_{1})-w_{2}\|_{|\pi|}\leq 2^{-2}$ and
$\|w_{2}\|_{\otimes_{n,s}E_{2}\,\widehat{\otimes}_{|\pi|}F_{2}}\leq\left(1+\frac{1}{2^{2}}\right)^{n}\|w_{2}\|_{|\pi|}+2^{n}\frac{1}{2^{2}}\leq 2^{n}\|w_{2}\|_{|\pi|}+2^{n-2}.$
Proceeding inductively, we may construct sequences $(E_{k})_{k}$ and $(F_{k})_{k}$ of finite dimensional vectors sublattices of $E$ and $F$, respectively, and a sequence $(w_{k})_{k}\subset\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$ such that $w_{k}\in\otimes_{n,s}E_{k}\,\otimes F_{k}$, $\|(w-w_{1}-\cdots-w_{k-1})-w_{k}\|_{|\pi|}\leq 2^{-k}$ and
$\|w_{k}\|_{\otimes_{n,s}E_{k}\,\widehat{\otimes}_{|\pi|}F_{k}}\leq 2^{n}\|w_{k}\|_{|\pi|}+2^{n-k}\,\text{ for every $k>1$.}$
Thus,
$\left|\|w\|_{|\pi|}-\sum_{i=1}^{k}\|w_{i}\|_{|\pi|}\right|\leq\left\|w-\sum_{i=1}^{k}w_{i}\right\|_{|\pi|}\leq 2^{-k}$
holds for every $k\in\mathbb{N}$. By letting $k\to\infty$, we get that $\|w\|_{|\pi|}=\sum_{i=1}^{\infty}\|w_{i}\|_{|\pi|}<\infty$ and $w=\displaystyle\sum_{k=1}^{\infty}w_{k}$. We also note that $\|w_{k+1}\|_{|\pi|}\leq\|w\|_{|\pi|}-\sum_{i=1}^{k}\|w_{i}\|_{|\pi|}\leq 2^{-k}$ holds for every $k\in\mathbb{N}$, and so
$\|w_{k}\|_{\otimes_{n,s}E_{k}\,\widehat{\otimes}_{|\pi|}F_{k}}$ $\leq 2^{n}\|w_{k}\|_{|\pi|}+2^{n-k}\leq 2^{n}\,2^{-(k-1)}+2^{n-k}$
$=2^{n}\,2^{-k+1}+2^{n-k}=2^{n-k}(2+1)=\frac{2^{n}\cdot 3}{2^{k}}$
for every $k\geq 2.$ Assuming $\|w\|_{|\pi|}=1$, for $k=1$, we have that
$\|w_{1}\|_{\otimes_{n,s}E^{\ast}\,\widehat{\otimes}_{|\pi|}F_{1}}\leq 2^{n}\|w_{1}\|_{|\pi|}+2^{n-1}\leq 2^{n}\left(\|w\|_{|\pi|}+\frac{1}{2}\right)\leq\displaystyle\frac{2^{n}\cdot 3}{2^{1}},$
and we are done. ∎
It was proved in [4] that, whenever $F$ is a Dedekind complete Banach lattice with the LAP, then $\overline{\Phi}\colon E^{\ast}\,\widehat{\otimes}_{|\pi|}F\to\mathcal{N}^{r}(E;F)$ is one-to-one, and hence an isometry, for every Banach lattice $E$. The same argument used in Blanco’s proof can be adapt to obtain a polynomial version of this result that will be presented in our next theorem. First, we recall that $\mathcal{P}_{f}(^{n}E;F)$ denotes the vector space of all finite rank $n$-homogeneous polynomials from $E$ into $F$.
Theorem 3.3.
Let $E$ and $F$ be Banach lattices with $F$ Dedekind complete. If $E^{*}$ and $F$ have the LAP, then the map $\overline{\Phi}_{n}\colon\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F\to\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;F)$ is injective, and hence an isometry.
Proof.
Suppose for the sake of contradiction that $\overline{\Phi}_{n}$ is not injective. Thus, there exists $w\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$ with $\|w\|_{|\pi|}=1$ such that $\overline{\Phi}_{n}(w)=0$. Since $E^{*}$ and $F$ are Dedekind complete, we can apply Lemma 3.2 to obtain sequences $(G_{k})_{k}$ and $(F_{k})_{k}$ of finite-dimensional vector sublattices of $E^{*}$ and $F$, respectively, and the corresponding sequence $(w_{k})_{k}\subset\widehat{\otimes}_{n,s,|\pi|}E^{*}\widehat{\otimes}_{|\pi|}F$. Since, for every $k\in\mathbb{N}$, $w_{k}\in\otimes_{n,s}G_{k}\otimes F_{k}$ and $\|w_{k}\|_{\otimes_{n,s}G_{k}\,\widehat{\otimes}_{|\pi|}F_{k}}\leq 2^{n}\|w_{k}\|_{|\pi|}+2^{n-k}\leq\displaystyle\frac{2^{n}\cdot 3}{2^{k}}$, there are a positive elements $z_{k,1},\dots,z_{k,m_{k}}\in\otimes_{n,s}G_{k}$ and positive vectors $y_{k,1},\dots,y_{k,m_{k}}\in F_{k}$ such that
$|w_{k}|\leq\displaystyle\sum_{j=1}^{m_{k}}z_{k,j}\otimes y_{k,j}\quad\text{ and}\quad\displaystyle\sum_{j=1}^{m_{k}}\|z_{k,j}\|_{s,|\pi|}\|y_{k,j}\|\leq\frac{2^{n}\cdot 3}{2^{k}}.$
Claim 1: For each $k\in\mathbb{N}$, there are $0\leq u_{k,1},\dots,u_{k,m_{k}}$ $\in$ $\otimes_{n,s}G_{k}$ and $0\leq v_{k,1},\dots,v_{k,m_{k}}\in F_{k}$ such that
$|w_{k}|\leq\sum_{j=1}^{m_{k}}u_{k,j}\otimes v_{k,j},\,\max_{1\leq j\leq m_{k}}\|v_{k,j}\|=\frac{1}{k},\,\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\|u_{k,j}\|_{s,|\pi|}<\infty.$ (1)
Indeed, first note that we can assume that $y_{k,j}\not=0$ as if it is equal to 0, then $u_{k,j}\otimes z_{k,j}=0$. For each $k\in\mathbb{N}$ and each $j=1,\dots,m_{k}$, we let
$u_{k,j}=k\|y_{k,j}\|z_{k,j}\quad\text{and}\quad v_{k,j}=\frac{y_{k,j}}{k\|y_{k,j}\|}.$
Thus, $0\leq u_{k,1},\dots,u_{k,m_{k}}$ $\in$ $\otimes_{n,s}G_{k}$ and $0\leq y_{k,1},\dots,y_{k,m_{k}}\in F_{k}$ satisfy
$|w_{k}|\leq\sum_{j=1}^{m_{k}}z_{k,j}\otimes y_{k,j}=\sum_{j=1}^{m_{k}}u_{k,j}\otimes v_{k,j},\quad\max_{1\leq j\leq m_{k}}\|v_{k,j}\|=\frac{1}{k},$
and
$\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\|u_{k,j}\|_{s,|\pi|}=\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}k\|z_{k,j}\|_{s,|\pi|}\|y_{k,j}\|\leq\sum_{k=1}^{\infty}\frac{2^{n}\cdot 3k}{2^{k}}<\infty.$
Claim 2: For each $k\in\mathbb{N}$ and each $j=1,\dots,m_{k}$, there are $g_{k,j}^{1},\dots,g_{k,j}^{p_{k,j}}\in G_{k}^{+}$ and $\widetilde{v}_{k,j}^{1},\dots,\widetilde{v}_{k,j}^{p_{k,j}}\in F_{k}^{+}$ such that $|w_{k}|\leq\displaystyle\sum_{j=1}^{m_{k}}\sum_{i=1}^{p_{k,j}}\theta_{n}(g_{k,j}^{i})\otimes\widetilde{v}_{k,j}^{i},$
$\max_{1\leq j\leq m_{k}}\max_{1\leq i\leq p_{k,j}}\|g_{k,j}^{i}\|=\frac{1}{k}\,\text{ and }\,\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\sum_{i=1}^{p_{k,j}}\|\widetilde{v}_{k,j}^{i}\|<\infty.$ (2)
First, for each $k\in\mathbb{N}$ and each $j=1,\dots,m_{k}$, there are $\varphi_{k,j}^{1},\dots,\varphi_{k,j}^{p_{k,j}}\in G_{k}^{+}$ such that $z_{k,j}\leq\sum_{i=1}^{p_{k,j}}\theta_{n}(\varphi_{k,j}^{i})$ and
$\sum_{i=1}^{p_{k,j}}\|\varphi_{k,j}\|^{n}\leq\|z_{k,j}\|_{s,|\pi|}+\frac{1}{m_{k}\displaystyle\max_{1\leq j\leq m_{k}}\|y_{k,j}\|\,2^{k}}.$
Next, we note that, as in Claim 1, we can assume that $\varphi_{k,j}\not=0$, and so we define
$g_{k,j}^{i}=\frac{\varphi_{k,j}^{i}}{k\|\varphi_{k,j}^{i}\|}\quad\text{and}\quad\widetilde{v}_{k,j}^{i}:=(k\|\varphi_{k,j}^{i}\|)^{n}y_{k,j}.$
$0\leq g_{k,j}^{i}\in G_{k}$ and $0\leq\widetilde{v}_{k,1}^{i},\dots,\widetilde{v}_{k,m_{k}}^{i}\in F_{k}$ satisfy
$|w_{k}|\leq\sum_{j=1}^{m_{k}}z_{k,j}\otimes y_{k,j}\leq\sum_{j=1}^{m_{k}}\sum_{i=1}^{p_{k,j}}\theta_{n}(g_{k,j}^{i})\otimes{\tilde{v}}^{i}_{k,j},\\ \quad\max_{1\leq j\leq m_{k}}\max_{1\leq i\leq p_{k,j}}\|g_{k,j}^{i}\|=\frac{1}{k},$
and
$\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\sum_{i=1}^{p_{k,j}}\|\widetilde{v}_{k,j}^{i}\|$ $=\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\sum_{i=1}^{p_{k,j}}k^{n}\|\varphi_{k,j}^{i}\|^{n}\|y_{k,j}\|$
$\leq\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}k^{n}\left(\|z_{k,j}\|_{s,|\pi|}+\frac{1}{m_{k}\max_{1\leq j\leq m_{k}}\|y_{k,j}\|2^{k}}\right)\|y_{k,j}\|$
$\leq\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}k^{n}\|z_{k,j}\|_{s,|\pi|}\|y_{k,j}\|+\sum_{k=1}^{\infty}\frac{k^{n}}{2^{k}}$
$\leq\sum_{k=1}^{\infty}\frac{2^{n}\cdot 3k^{n}}{2^{k}}+\sum_{k=1}^{\infty}\frac{k^{n}}{2^{k}}<\infty.$
For each $k,j,i$, define
$G_{k,j}^{i}:=\left\{{g\in G_{k}}\,:\,{|g|\leq mg_{k,j}^{i}\,\text{ for some }m\in\mathbb{N}}\right\}$
and
$F_{k,j}=\left\{{y\in F_{k}}\,:\,{|y|\leq m\,v_{k,j}\,\text{ for some }m\in\mathbb{N}}\right\}.$
Ordering $(v_{k,j})$ by
$v_{1,1},\dots,v_{1,m_{1}},v_{2,1}\dots,v_{2,m_{2}},\dots$
and $(g_{k,j}^{i})$ by
$g_{1,1}^{1},\dots,g_{1,1}^{p_{1}},g_{1,2}^{1},\dots,g_{1,2}^{p_{1,2}},\dots,g_{1,m_{1}}^{1},\dots,g_{1,m_{1}}^{p_{1,m_{1}}},\dots$
we get that $(G_{k,j}^{i},g_{k,j}^{i})$ and $((F_{k,j}),v_{k,j})$ are sequences of principal pairs of $E^{*}$ and $F$, respectively. Let $\varepsilon>0$. As $E^{*}$ and $F$ have the LAP, there are finite rank operators $S\colon E^{*}\to E^{*}$ and $T\colon F\to F$ such that
$\||(S-id)_{G_{k,j}^{i}}|(g_{k,j}^{i})\|<\varepsilon\,\text{ and }\,\||(T-id)_{F_{k,j}}|(v_{k,j})\|<\varepsilon$
for all $k,j,i$.
Defining $\Psi\colon\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F\to\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$ by $\Psi(\theta_{n}(x^{*})\otimes y)=\theta_{n}(S(x^{*}))\otimes T(y)$ for all $x^{*}\in E^{\ast}$ and $y\in F$, and extending by linearity and continuity, we get that $\Psi(w_{k})\in\otimes_{n,s}\,S(E^{*})\otimes T(F)$ for every $k\in\mathbb{N}$. To see that $\Psi$ is well-defined we note that if $\sum_{i=1}^{k}\theta_{n}(x_{i}^{*})\otimes y_{i}=\sum_{j=1}^{l}\theta_{n}(z_{j}^{*})\otimes w_{j}$, then
$\sum_{i=1}^{k}\theta_{n}(Sx_{i}^{*})\otimes Ty_{i}$ $=\sum_{i=1}^{k}\varphi_{1}^{\otimes}(\theta_{n}(x_{i}^{*}))\otimes Ty_{i}=\varphi_{1}^{\otimes}\otimes T(\sum_{i=1}^{k}\theta_{n}(x_{i}^{*})\otimes y_{i})$
$=\varphi_{1}^{\otimes}\otimes T(\sum_{j=1}^{l}\theta_{n}(z_{j}^{*})\otimes w_{j})=\sum_{j=1}^{l}\varphi_{1}^{\otimes}(\theta_{n}(z_{j}^{*}))\otimes Tw_{j}$
$=\sum_{j=1}^{l}\theta_{n}(Sz_{j}^{*})\otimes Tw_{j},$
where $\varphi_{1}^{\otimes}$ is the linearization of the regular symmetric $n$-linear map $\varphi_{1}:E^{*}\times\cdots\times E^{*}\to\widehat{\otimes}_{n,s,|\pi|}E^{*}$ given by $\varphi_{1}(x_{1}^{*},\dots,x_{n}^{*})=Sx_{1}^{*}\otimes_{s}\cdots\otimes_{s}Sx_{n}^{*}$.
Moreover, as $\otimes_{n,s}S(E^{*})\otimes T(F)$ is closed in $\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}F$ and $\Psi(w)=\displaystyle\sum_{k=1}^{\infty}\Psi(w_{k})$, we have that $\Psi(w)\in\otimes_{n,s}S(E^{*})\otimes T(F)$, which yields that $\overline{\Phi}_{n}(\Psi(w))=\Phi_{n}(\Psi(w))$, where $\Phi_{n}$ is the following isomorphism of vector spaces:
$\Phi_{n}\colon\otimes_{n,s}E^{\ast}\otimes F\to\mathcal{P}_{f}(^{n}E;F),\quad\Phi_{n}(\theta_{n}(x^{*})\otimes y)=(x^{*}(x))^{n}y.$
Now, considering $\Gamma:\mathcal{P}_{\cal N}(^{n}E;F)\to\mathcal{P}_{\cal N}(^{n}E;F)$ given by
$\Gamma\left(\sum_{i=1}^{\infty}(x_{i}^{*}(\cdot))^{n}y_{i}\right)=\sum_{i=1}^{\infty}(Sx_{i}^{*}(\cdot))^{n}Ty_{i},$
we have that
$\Phi_{n}(\Psi(\theta_{n}(x^{*})\otimes y))=\Phi_{n}(\theta_{n}(Sx^{*})\otimes Ty)=(Sx^{*}(\cdot))^{n}Ty=\Gamma((x^{*}(\cdot))^{n}y)=\Gamma(\Phi_{n}(\theta_{n}(x^{*})\otimes y))$
holds for all $x^{*}\in E^{*}$ and $y\in F$, hence $\Phi_{n}(\Psi(w_{k}))=\Gamma(\Phi_{n}(w_{k}))$ for every $k\in\mathbb{N}$. Therefore
$\Phi_{n}(\Psi(w))$ $=\overline{\Phi}_{n}(\Psi(w))=\sum_{k=1}^{\infty}\overline{\Phi}_{n}(\Psi(w_{k}))=\sum_{k=1}^{\infty}\Phi_{n}(\Psi(w_{k}))$
$=\sum_{k=1}^{\infty}\Gamma(\Phi_{n}(w_{k}))=\Gamma\left(\sum_{k=1}^{\infty}\Phi_{n}(w_{k})\right)=\Gamma\left(\sum_{k=1}^{n}\overline{\Phi}_{n}(w_{k})\right)=\Gamma(\overline{\Phi}_{n}(w))=0.$
Now, since $\Phi_{n}$ is a linear isomorphism, we obtain that $\Psi(w)=0$. From this, we will prove that $w=0$. First, we notice that, for each $k\in\mathbb{N}$
$|(id-\Psi)(w_{k})|$ $\leq|(id-id\otimes T)(w_{k})|+|(id\otimes T-\Psi)(w_{k})|$
$=|id\otimes(id_{F}-T)(w_{k})|+|(id-\otimes^{n}S)\otimes T(w_{k})|$
$\leq|id\otimes(id_{F}-T)|(|w_{k}|)+|(id-\otimes^{n}S)\otimes T|(|w_{k}|)$
$\leq|id\otimes(id_{F}-T)|\left(\sum_{j=1}^{m_{k}}u_{k,j}\otimes v_{k,j}\right)+$
$+|(id-\otimes^{n}S)\otimes T|\left(\sum_{j=1}^{m_{k}}\sum_{i=1}^{p_{k,j}}\theta_{n}(g_{k,j}^{i})\otimes\widetilde{v}_{k,j}^{i}\right),$
where in the last inequality we used the constructions obtained in Claim 1 and Claim 2 above. We will divide our proof in two steps now. First:
$|id\otimes(id_{F}-T)|\left(\sum_{j=1}^{m_{k}}u_{k,j}\otimes v_{k,j}\right)$ $=\sum_{j=1}^{m_{k}}|id\otimes(id_{F}-T)|(u_{k,j}\otimes v_{k,j})$
$=\sum_{j=1}^{m_{k}}u_{k,j}\otimes|(id-T)_{F_{k,j}}|(v_{k,j})$
Second:
$|(id-\otimes^{n}S)\otimes T|$ $\left(\sum_{j=1}^{m_{k}}\sum_{i=1}^{p_{k,j}}\theta_{n}(g_{k,j}^{i})\otimes\widetilde{v}_{k,j}^{i}\right)=\sum_{j=1}^{m_{k}}\sum_{i=1}^{p_{k,j}}|(id-\otimes^{n}S)\otimes T|(\theta_{n}(g_{k,j}^{i})\otimes\widetilde{v}_{k,j}^{i})$
$=\sum_{j=1}^{m_{k}}\sum_{i=1}^{p_{k,j}}\theta_{n}(|(id-S)_{G_{k,j}^{i}}|(g_{k,j}^{i}))\otimes T(\widetilde{v}_{k,j}^{i})$
Combining the above, we obtain
$\|w\|_{|\pi|}$ $=\|w-\Psi(w)\|_{|\pi|}=\|(id-\Psi)(\sum_{k=1}^{\infty}w_{k})\|_{|\pi|}\leq\sum_{k=1}^{\infty}\|(id-\Psi)(w_{k})\|_{|\pi|}$
$\leq\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\|u_{k,j}\|\||(id-T)_{F_{k,j}}|(v_{k,j})\|+\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\sum_{i=1}^{p_{k,j}}\||(id-S)_{G_{k,j}^{i}}|(g_{k,j}^{i})\|^{n}\|T(\widetilde{v}_{k,j}^{i})\|$
$\leq\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\|u_{k,j}\|\varepsilon+\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\sum_{i=1}^{p_{k,j}}\varepsilon^{n}\|T\|\|\widetilde{v}_{k,j}^{i}\|.$
Taking $\varepsilon\longrightarrow 0$, we obtain $\|w\|_{|\pi|}=0$, and we have a contradiction. ∎
Corollary 3.4.
Let $E,F,G$ be Banach lattices such that $G$ is Dedekind complete and $E^{*},F^{*}$ and $G$ have the LAP. If $E^{*}$ and $F^{*}$ are lattice isomorphic, then $\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;G)$ and $\mathcal{P}_{\mathcal{N}}^{r}(^{n}F;G)$ are lattice isomorphic.
Proof.
If $E^{*}$ and $F^{*}$ are lattice isomorphic, then $\widehat{\otimes}_{n,s,|\pi|}E^{\ast}$ and $\widehat{\otimes}_{n,s,|\pi|}F^{\ast}$ are lattice isomorphic. Indeed, if $\varphi\colon E^{*}\to F^{*}$ is a lattice isomorphism, then $P(x^{*})=\theta_{n}(\varphi(x^{*}))$ defines a positive $n$-homogeneous polynomial from $E$ into $\widehat{\otimes}_{n,s,|\pi|}F^{\ast}$, and so $P^{\otimes}\colon\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\to\widehat{\otimes}_{n,s,|\pi|}F^{\ast}$ is a positive linear operator such that $P^{\otimes}(\theta_{n}(x^{*}))=\theta_{n}(\varphi(x^{*}))$ for every $x^{*}\in E^{*}$. Analogously, there exists a positive operator $S\colon\widehat{\otimes}_{n,s,|\pi|}F^{\ast}\to\widehat{\otimes}_{n,s,|\pi|}E^{\ast}$ such that $S(y^{*})=\theta_{n}(\varphi^{-1}(y^{*}))$ for every $y^{*}\in F^{*}$. Finally, since
$S\circ P^{\otimes}(\theta_{n}(x^{*}))=S(\theta_{n}(\varphi(x^{*})))=\theta_{n}(\varphi^{-1}(\varphi(x^{*})))=\theta_{n}(x^{*})$
holds for every $x^{*}\in E^{*}$, we obtain that $S\circ P^{\otimes}(z)=z$ for every $z\in\widehat{\otimes}_{n,s,|\pi|}E^{\ast}$. The same idea also proves that $P^{\otimes}\circ S(z)=z$ for every $z\in\widehat{\otimes}_{n,s,|\pi|}F^{\ast}$. Thus, $(P^{\otimes})^{-1}=S$, and by [1, Theorem 2.15] $P^{\otimes}$ is a lattice isomorphism. An adaptation of the above argument can be used to prove that $\widehat{\otimes}_{n,s,|\pi|}E^{\ast}\,\widehat{\otimes}_{|\pi|}G$ is lattice isomorphic to $\widehat{\otimes}_{n,s,|\pi|}F^{\ast}\,\widehat{\otimes}_{|\pi|}G$ . Now, by applying Theorem 3.3, we get that $\mathcal{P}_{\mathcal{N}}^{r}(^{n}E;G)$ and $\mathcal{P}_{\mathcal{N}}^{r}(^{n}F;G)$ are lattice isomorphic. ∎
We conclude the article by providing an example where Corollary 3.4 can be applied:
Example 3.5.
It follows from [12, Remark 4.3] that every atomic Banach lattice with an order continuous norm has the LAP. So, letting $E$ and $F$ be $c_{0},c$ or the AM-space $X$ from Example 2.6, we obtain that $E^{*}$ and $F^{*}$ are both lattice isomorphic to $\ell_{1}$, hence $E^{*}$ and $F^{*}$ are lattice isomorphic with the LAP. Now, if $G$ is an arbitrary atomic Banach lattice with order continuous norm, we conclude by Corollary 3.4 that $\mathcal{P}^{r}_{\cal N}(^{n}E;G)$ and $\mathcal{P}^{r}_{\cal N}(^{n}F;G)$ are lattice isomorphic.
C. Boyd
Department of Mathematics and Statistics
University College Dublin
Belfield – Dublin 4 – Ireland
e-mail: christopher.boyd@ucd.ie
V. C. C. Miranda
Departamento de Matemática
Instituto de Ciências Matemáticas e de Computação
Universidade de São Paulo
13566-590 – São Carlos - SP – Brazil
e-mail: viniciusmiranda@icmc.usp.br